### v1

# Two-sample comparison through additive tree models for density ratios

Abstract
The ratio of two densities characterizes their differences. We consider learning the density ratio given i.i.d. observations from each of the two distributions. We propose additive tree models for the density ratio along with efficient algorithms for training these models using a new loss function called the balancing loss. With this loss, additive tree models for the density ratio can be trained using algorithms original designed for supervised learning. Specifically, they can be trained from both an optimization perspective that parallels tree boosting and from a (generalized) Bayesian perspective that parallels Bayesian additive regression trees (BART). For the former, we present two boosting algorithms—one based on forward-stagewise fitting and the other based on gradient boosting, both of which produce a point estimate for the density ratio function. For the latter, we show that due to the loss function’s resemblance to an exponential family kernel, the new loss can serve as a pseudo-likelihood for which conjugate priors exist, thereby enabling effective generalized Bayesian inference on the density ratio using backfitting samplers designed for BART. The resulting uncertainty quantification on the inferred density ratio is critical for applications involving high-dimensional and complex distributions in which uncertainty given limited data can often be substantial. We provide insights on the balancing loss through its close connection to the exponential loss in binary classification and to the variational form of $f$-divergence, in particular that of the squared Hellinger distance. Our numerical experiments demonstrate the accuracy and computational efficiency of the proposed approach while providing unique capabilities in uncertainty quantification. We demonstrate the application of our method in a case study involving assessing the quality of generative models for microbiome compositional data.
1 Introduction
A vast range of applications involve comparing two groups of observations to assess their differences. Classical formulation of two-sample comparison often involves parametric or semiparametric assumptions on the underlying distributions and focuses on testing the null hypothesis of no difference with or without a given alternative in mind. However, this hypothesis testing perspective is often inadequate in most modern applications of two-sample comparison. The practitioners nowadays typically also desire to learn the specific nature of the underlying difference. This problem lies at the core of numerous other modern statistical analyses and machine learning objectives.
For example, large-scale biomedical studies involving comparing genetic measurements and various biomarkers of patients and controls, there is often no question that some types of two-sample differences do exist, and with a sufficiently large sample size, the null of no difference at all will be rejected. It is more helpful if one can specifically pinpoint what such differences are to help understand their biological meaning. Another timely application is the assessment of generative models. The key to assessing the quality of different models and algorithms for generating synthetic samples that imitate the real sampling distribution of the training data lies in deciphering the distributional difference between the real sampling distribution and that induced by the generative models, which is essentially a two-sample comparison problem. Some generative models, such as generative adversarial networks (GANs) (Goodfellow et al., 2014), in essence iteratively apply two-sample comparison and use the estimated misfit between the two distributions to sequentially improve the generative model. Two-sample comparison is also a quintessential problem in causal inference, in which consistent estimation of treatment effects often relies on correctly adjusting for the difference in the covariate distributions between the treated and the control groups.
Our focus here is on two-sample comparison in a nonparametric fashion using the density ratio function as a functional “statistic” for characterizing two-sample differences. The density ratio fully summarizes how two distributions differ across the sample space, and thus allows one to not only carry out hypothesis testing but to identify the nature of the difference as well. If, for example, the estimated density ratio “significantly” deviates away from 1 on parts of the sample space, it will allow us to draw statistical conclusions regarding whether, where, and how the two distributions differ. At the same time, the notion of “significant” deviation will require some form of uncertainty quantification on the inferred density ratio.
There is a general misconception that nonparametric density ratio learning is “too hard” for complex, high-dimensional data because density estimation itself in such settings is already a hard enough problem in multivariate. Why then would one attempt an even harder problem of learning the ratio of two densities? Though it might appear counter-intuitive, we argue that density ratio estimation (DRE) is in fact a much easier problem than density estimation in most modern applications. A density by definition is always a density ratio, with the denominator being the corresponding base measure against which the density is computed. The typical base measure, for example the Lebesgue measure or the uniform density in continuous cases, is generally chosen out of mathematical convenience and the lack of knowledge that takes no account of the dataset at hand. It generally does not resemble any sampling distribution in high-dimensional problems. When good reliable knowledge of the sampling distribution does exist, one would be in the classical setting of a “goodness-of-fit” problem. In contrast, in most applications of two-sample comparison, the two distributions are often quite similar to each other and can serve as a “baseline of reference” for each other. As such, the density ratio between these two samples then is often much simpler than the underlying densities (i.e., their ratios with respect to the Lebesgue measure).
This is the fundamental reason behind a well-known observation in the DRE literature, that though estimating a density ratio can be done by first separately estimating the two densities and then computing their ratios based on the density estimates, it is usually more effective to adopt a method directly aimed at the density ratio, especially when the difference exists only in a small subset of the sample space or in a marginal distribution. Several authors have adopted this strategy. Some propose kernel-based methods (e.g. Sugiyama et al., 2007b; Kanamori et al., 2009; Hido et al., 2011; Liu et al., 2013; Yamada et al., 2013), and various kinds of other methods are proposed, such as an estimation via classification (Rhodes et al., 2020; Choi et al., 2021, 2022) and a deep neural network method (Kato and Teshima, 2021).
Beyond the stand-alone two-sample comparison problem which is our focus herein, DRE has also been applied as a component of various types of learning methods in which two-sample comparison arises as a sub-problem directly or indirectly. A far-from-exhaustive list includes an estimation of the mutual information (Suzuki et al., 2009a, b), conditional density estimation (Sugiyama et al., 2010), covariate shift correlation (Shimodaira, 2000; Sugiyama et al., 2007a; Reddi et al., 2015; Stojanov et al., 2019), outlier detection (Hido et al., 2011; Kato and Teshima, 2021), causal inference (Uehara et al., 2020), and change point detection (Kawahara and Sugiyama, 2009; Liu et al., 2013; du Plessis et al., 2015). Additionally, estimating density ratios is utilized in Markov Chain Monte Carlo (MCMC) for intractable likelihoods (Kaji and Ročková, 2023).
We list several desiderata in learning the density ratio and discuss how we aim to achieve them. First, the model or function class representing the density ratio should be flexible enough to effectively approximate complex forms of the underlying truth. It should scale well in both sample size and in dimensionality. To these ends, we propose a class of density ratio learners based on additive tree boosting (see e.g. Friedman et al., 2000; Hastie et al., 2009). Additive tree ensembles are a natural candidate for the density ratio problem, because of their success in classification problems. In fact, in the machine learning literature, a widely used technique to learn the density ratio between two samples is the so-called “density-ratio trick” (Sugiyama et al., 2012; Durkan et al., 2020; Thomas et al., 2022), which involves first fitting a binary classifier to the two samples, and then, based on the estimated posterior odds of the sample label, compute the density ratio. Additive tree boosting such as AdaBoost is among the most effective for binary classification on the so-called “tabular data”. We show that DRE does not have to involve the indirect route of inverting a binary classifier based on additive tree ensembles, but can directly be learned from the latter. As we will show, the resulting estimate will often be substantially more accurate than what the density ratio trick will produce, especially when the sample size is unbalanced or the two-sample difference is highly local; in both of these scenarios, losses designed for binary classification are a poor proxy for learning the density ratio.
We introduce two tree boosting algorithms for estimating the density ratio. We use an additive tree ensemble as the function class for approximating the density ratio. by adding tree-based weak learners. The first algorithm is based on forward-stagewise fitting (Hastie et al., 2009) and it operationally resembles Adaboost. In each iteration the observations are weighted according to the discrepancies between the current estimate and the actual responses, so that in each step we can fit a new weak learner focusing on observations with large discrepancies. This algorithm sequentially and greedily minimizes the balancing loss through fitting a single-tree piecewise constant approximation to the density ratio of the weighted samples. The second algorithm is under the general paradigm of gradient boosting (Friedman, 2001).
These algorithms inherit important advantages of supervised tree boosting — they can be implemented with a small computation cost, and at the same time achieve good numerical performance, especially when the difference lies in a marginal distribution of a subset of the variables. One can incorporate regularization to avoid overfitting effectively by making the contribution of each base learner weak with a small learning rate, following the standard approach in supervised boosting (Hastie et al., 2009; Friedman, 2001).
The second desideratum is to attain uncertainty quantification for the estimated density ratio, thereby allowing statistical inference to proceed. This aspect has mostly been ignored in the existing literature on DRE. One exception was Bayesian nonparametric models for the DR using a single tree model (Ma and Wong, 2011; Soriano and Ma, 2017). While they provide means to examine posterior uncertainty in a fully Bayesian manner, piecewise-constant functions represented by single trees are very restrictive and therefore inadequate for approximating complex density ratios. Additive trees substantially enrich the class of functions, but they do not allow a simple likelihood formulation for modeling density ratios, and consequently, fully Bayesian uncertainty quantification is challenging. To overcome this, we propose a generalized Bayesian approach Bissiri et al. (2016) using the balancing loss to construct a pseudo-likelihood for this model class that has shown excellent performance in supervised problems, as realized by Bayesian additive and regression trees (BART) (Chipman et al., 2010). Interestingly, we show that due to the resemblance of the balancing loss to an exponential family kernel, there exists a natural conjugate prior for the model parameters under the pseudo-likelihood which allows standard sampling strategies for BART models (Chipman et al., 2010) to be directly adopted for sampling the Gibbs posterior on the density ratio.
The remainder of the paper is organized as follows. In Section 2 we start by introducing the balancing loss for learning the density ratio and draw its connection to the variational form of squared Hellinger distance and to the exponential loss for binary classification. We then introduce an additive tree model on the (log) density ratio along with two boosting algorithms for fitting this model based on forward stage-wise fitting and gradient boosting, respectively. In Section 3, we introduce a generalized Bayesian inference framework that uses the balancing loss as a (log) pseudo-likelihood and show one can accomplish Bayesian backfitting sampling on the model using standard algorithms for BART through exploiting the analogy of the pseudo-likelihood to an exponential family. In Section 4, we evaluate the performance of the proposed methods in comparison with the existing approaches in lower and higher-dimensional simulation settings followed by a case study involving assessing and comparing multiple generative models for microbiome compositions in Section 5. We give a few concluding remarks in Section 6.
2 Method
2.1 The balancing loss for density ratio learning
Let $P$ and $Q$ be unknown probability measures defined on the common sample space denoted by $\Omega$. We focus on the case where $\Omega$ is a $d$-dimensional Euclidean space. We assume that the two measures $P$ and $Q$ admit density functions with respect to the Lebesgue measure $\mu$ and they are denoted by $p$ and $q$, respectively. Our goal is to estimate the unknown density ratio $p/q$ based on two i.i.d. samples.
Let $r$ denote an estimate of the density ratio. In the following discussion, we assume that $r$ is positive $\mu$-almost everywhere. To introduce an algorithm for optimizing $r$, we define a loss function called the balancing loss based on the following fact: If $r$ is equal to the true density ratio $p/q$, then for any measurable subset $B\subset\Omega$, we have
$\int_{B}r^{-1/2}pd\mu=\int_{B}r^{1/2}qd\mu=\int_{B}\sqrt{pq}d\mu.$ (1)
When $B=\Omega$, the whole sample space, this corresponds to the affinity between $p$ and $q$. The first equation implies that if we weight the two samples respectively with $r^{-1/2}$ and $r^{1/2}$, their probability over any set $B$ becomes equal. In this sense, we say the two samples are “balanced”. Motivated by this observation, we introduce the notation $w=r^{1/2}$ to denote the balancing function, which is intended to approximate the square root of the density ratio $\sqrt{p/q}$. We then consider the following balancing loss function:
$l(w)=\mathbb{E}_{p}[w^{-1}]+\mathbb{E}_{q}[w].$
In practice, we observe i.i.d. draws from $p$ and $q$:
$x^{(0)}_{1},\dots,x^{(0)}_{n_{0}}\sim p,\ x^{(1)}_{1},\dots,x^{(1)}_{n_{1}}\sim q.$
The finite-sample version of the balancing loss is defined by
$l_{n}(w)=\frac{1}{n_{0}}\sum^{n_{0}}_{i=1}w^{-1}(x^{(0)}_{i})+\frac{1}{n_{1}}\sum^{n_{1}}_{i=1}w(x^{(1)}_{i})$,
where $n=n^{(0)}+n^{(1)}$ indicates the total number of observations.
2.2 Connections to existing concepts in statistical learning
While the design of the balancing loss may appear ad hoc at first, next we show that it can be naturally motivated from two well-established perspectives in machine learning.
2.2.1 Connection 1: the “density ratio trick” based on binary classification
The first perspective is related to a common technique widely adopted in the machine learning community for estimating the density, often referred to as the “density ratio trick”, which involves turning the problem into a supervised, binary classification on the two group labels (Sugiyama et al., 2012; Durkan et al., 2020; Thomas et al., 2022). It involves inverting a binary classifier to compute the density ratio. We show that the balancing loss arises naturally in this context as a variant of the exponential loss widely used in this context, but only the balancing loss targets the density ratio directly and thus is more robust than the density ratio trick when the two sample sizes are unbalanced and/or the two-sample differences exist locally in a small subset of the sample space. These will also be validated experimentally later in our numerical examples where a direct comparison to the density ratio trick will be provided.
The balancing loss can be seen as a modification of the exponential loss, which is used for classification with the Adaboost, described in Friedman et al. (2000). To see this, we suppose that in generating the two samples, one first randomly decides which sample to draw the observation from. That is, one first draws a random variable $y$ that takes 1 and -1 with a 50-50 chance. Given $y$ one generates the corresponding sample $x$ from $P$ if $y=1$ and from $Q$ if y=1y=-1, respectively. Then it is straightforward to show that
$\mathbb{E}_{p}[w^{-1}]+\mathbb{E}_{q}[w]\propto\mathbb{E}[e^{-yF}],$
where $F=\log w$ and the expectation in the latter is taken with respect to the joint distribution of $x$ and $y$. What we have on the right-hand side is exactly the exponential loss (Friedman et al., 2000; Hastie et al., 2009).
We note that the exponential loss used in AdaBoost in fact integrates $y$ with respect to its marginal distribution from the data. While the population minimizer for the AdaBoost is the log odds between the two samples (Friedman et al., 2000), which is proportional to the density ratio, one may wonder then why the original AdaBoost cannot suffice for learning density ratios since one can simply multiply a constant (i.e., the ratio of the proportion of the two samples) to get an estimate of the density ratio. The main reason is that by weighting the loss using the marginal proportion of two samples, vanilla AdaBoost prioritizes classification accuracy over density ratio estimation; this can be most clearly demonstrated when there are parts of the sample space in which one of the samples is much more abundant than the other, which can occur when the sample sizes are very different or when the two sample densities have different supports or tail behaviors. Such cases are in fact quite common in multivariate applications, and numerical evaluations will be provided in Section 4.
Connection 2: The variational form of the squared Hellinger distance
The design of the balancing loss can also be motivated from the perspective of estimating the $f$-divergence between $p$ and $q$. In particular, in this section we show that the balancing loss turns out to be exactly equivalent to the variational form of one particular $f$-divergence, namely the squared Hellinger distance.
For a given convex function $f$, the general $f$-divergence is defined as $D_{f}(P\|Q):=\int f(\frac{p}{q})dQ$. It is known (Rényi, 1961; Csiszár, 1964; Ali and Silvey, 1966; Nguyen et al., 2010; Polyanskiy and Wu, 2025) that the dual form can be written as an optimization problem with respect to a functional class $\mathcal{F}$ mapping from $\mathcal{X}$ to $\mathbb{R}$,
$D_{f}(P\|Q)\geq\sup_{g\in\mathcal{F}}\int[gdP-f^{c}(g)dQ]$ (2)
where $f^{c}$ is the conjugate dual function of $f$, defined as $f^{c}:=\sup_{u\in\mathbb{R}}\{uv-f(u)\}$. The dual objective is referred to as the variational form of the $f$-divergence. In the case of the squared Hellinger divergence, $f(u)=\frac{1}{2}(\sqrt{u}-1)^{2}$, and the dual objective is $\sup_{g<\frac{1}{2}}\Big{\{}\mathbb{E}_{P}[g(X)]\;-\;\mathbb{E}_{Q}\!\Big{[}\frac{g(X)}{1-2g(X)}\Big{]}\Big{\}}$ with the optimum value $g^{*}=\frac{1}{2}-\frac{1}{2}\sqrt{\frac{q}{p}}$, based on a direct application of Lemma 1 in Nguyen et al. (2010). Now let $w=(1-2g)^{-1}$, the variational form for the squared Hellinger divergence becomes
$\sup_{w>0}\left\{1-\frac{1}{2}\int w^{-1}pd\mu-\frac{1}{2}\int wqd\mu\right\}$ (3)
which is clearly equivalent to minimizing the balancing loss. In other words, learning the density ratio by minimizing the balancing loss is essentially estimating the squared Hellinger distance using its variational form. Based on the known minimizer of the variational form, the minimizer of the balancing loss is $w^{*}=\sqrt{p/q}$. Hence, learning the density ratio can be achieved through minimizing the balancing loss.
Beyond these two additional motivations of the balancing loss, we will see later that the symmetry of the balancing loss, in $w$ vs $w^{-1}$ and $p$ vs $q$ will become handy when designing a generalized Bayesian framework to carry out inference on additive tree models.
2.3 Choice of the function class: additive tree models
Given the close relationship between density ratios and posterior odds, the latter of which can be estimated using supervised tree boosting with additive tree models, it is natural to consider the class of additive trees as a promising candidate for approximating the density ratio. We thus consider from now on fitting additive tree models to the density ratio under the balancing loss.
We begin with a non-Bayesian approach that aims to find a (functional) point estimate of the density ratio by optimizing the finite-sample version of the balancing loss. The resulting training algorithms are very similar to supervised tree boosting. Before introducing the algorithms, we describe the additive tree model. We consider a model for the log-balancing weight $\log w$ in the form
$\log w=\sum_{k=1}^{K}f_{k}$ (4)
where each base learner $f_{k}$ is a piecewise constant function defined on the leaf partition of the sample space given a tree $T_{k}$ with leaf node parameters $\beta_{k}=\{\beta_{k}(A)\}_{A\in T_{k}}$ and written as
$f_{k}=\sum_{A\in T_{k}}\beta_{k}(A)\mathbf{1}_{A}.$ (5)
Under this model, we now are presented with the following optimization problem:
$\min_{\{f_{k}\}}l_{n}(w)=\min_{\{\beta_{k},T_{k}\}}\frac{1}{n_{0}}\sum^{n_{0}}_{i=1}w^{-1}(x^{(0)}_{i})+\frac{1}{n_{1}}\sum^{n_{1}}_{i=1}w(x^{(1)}_{i})$
Next, we introduce two algorithms for fitting the above model under the balancing loss. These are direct counterparts of two different approaches to fitting additive trees in supervised learning problems. The first corresponds to a forward-stagewise (FS) algorithm based on greedy steepest descent on the space of single tree-based piecewise-constant functions; the other corresponds to Friedman (2001)’s gradient boosting (GB) algorithm. In practice, these algorithms typically produce estimates that are very similar to each other, and so one could just adopt the simpler GB algorithm, but because the details of the FS algorithm shed more light on what constitutes effective function approximators for the density ratio under the balancing loss, we believe they are worthy of elaboration here.
Algorithm 1: A forward-stagewise algorithm
The following proposition characterizes the optimal single tree fit in each iteration of the boosting algorithm.
Proposition 2.1.
Given the current estimate $w_{k-1}$, the loss is minimized when the following conditions are satisfied:
  1. 1.
    For each $A\in T_{k}$,
    $e^{\beta_{k}(A)}=\sqrt{\frac{\int_{A}pw^{-1}_{k-1}d\mu}{\int_{A}qw_{k-1}d\mu}}.$
  2. 2.
    Let $P_{k}|_{T_{k}}$ and $Q_{k}|_{T_{k}}$ be discrete distributions with their probability masses indexed by $A\in T_{k}$ such that
    $P_{k}|_{T_{k}}(A)\propto\int_{A}w_{k-1}^{-1}pd\mu,\ Q_{k}|_{T_{k}}(A)\propto\int_{A}w_{k-1}qd\mu.$
    Then, $T_{k}$ is a partition structure that maximizes the Hellinger distance
    $H(P_{k}|_{T_{k}},Q_{k}|_{T_{k}})=\frac{1}{\sqrt{2}}\sqrt{\sum_{A\in T_{k}}\left(P_{k}|_{T_{k}}(A)-Q_{k}|_{T_{k}}(A)\right)^{2}.}$
The proof is provided in Supplementary Materials A.
Remark 1: If the balancing weight is correctly estimated, it achieves the symmetry between the two expectations in the loss in that $\mathbb{E}_{p}[w^{-1}]=\mathbb{E}_{q}[w]$, and we can impose this condition as a restriction on the estimator so that it does not focus on minimizing only one expectation in the two. As its justification, Proposition 2.1 implies that the balancing loss can be improved by multiplying the current estimate $w_{k}$ by a constant ($c$, say) such that $\mathbb{E}_{p}[(cw_{k})^{-1}]=\mathbb{E}_{q}[cw_{k}]$ (see Supplementary Materials A for the detail). Based on the discussion, we introduce this correction step in the boosting algorithms.
Remark 2: Given the connection of the balancing loss to the variational form of the Hellinger divergence, it is not surprising that the “optimal” single-tree in each iteration corresponds to the partition that maximizes the corresponding discretized approximation to the Hellinger divergence.
The optimal learner in the popular gradient-based boosting is obtained based on the first order condition in Friedman et al. (2000), while the proposition gives a solution that exactly maximizes the reduction in the balancing loss, or equivalently the increase in Hellinger distance. This is also equivalent to minimizing the discretized affinity, $\sum_{A\in T_{k}}\sqrt{P_{k}|_{T_{k}}(A)Q_{k}|_{T_{k}}(A)}.$
While in practice searching for the exact optimal tree in each step is not computationally tractable, this inspires a greedy coarse-to-fine algorithm in which we iteratively choose a single cut point that achieves the steepest ascent in Hellinger distance or the steepest ascent in the corresponding affinity. Specifically, when we consider splitting the current node $A$, we can search for the splitting rule such that for the children nodes $A_{l}$ and $A_{r}$ the following affinity is minimized:
$\sqrt{P_{k}|_{T_{k}}(A_{l})Q_{k}|_{T_{k}}(A_{l})}+\sqrt{P_{k}|_{T_{k}}(A_{r})Q_{k}|_{T_{k}}(A_{r})}.$
Algorithm 2: A gradient boosting algorithm
We can also introduce a gradient-based tree learner (Friedman, 2001). If a current estimate is $w$, with the log-transformation
$F=\log w,$
we formulate an ensemble where learners are combined additively. Then we can construct a tree partition that fits the functional gradients (i.e., residuals) defined for the two groups. As the two integrands are written as $e^{-F(x)}$ and $e^{F(x)}$, the negative gradients are as follows:
$\frac{\partial l_{n}(w)}{\partial F(x_{i}^{(0)})}=-\frac{1}{n_{0}}e^{-F(x_{i}^{(0)})}=\frac{1}{n_{0}}w^{-1}(x_{i}^{(0)})\quad\text{and}\quad-\frac{\partial l_{n}(w)}{\partial F(x_{i}^{(1)})}=\frac{1}{n_{1}}e^{F(x_{i}^{(1)})}=-\frac{1}{n_{1}}w(x_{i}^{(1)}).$
We follow the recipe provided in (Friedman, 2001): we first fit the new tree partition to the gradients by minimizing squared error loss, that is, minimizing variances of the gradients in the leaf nodes. To this end, we adopt the coarse-to-fine greedy search and for each node $A$, we select the splitting rule with children nodes $A_{l}$ and $A_{r}$ that minimizes the variances. After constructing the tree, secondly, we obtain the optimal values of the node parameters, which are already given by Proposition 2.1.
2.4 Regularization strategies
The success of tree boosting under the two algorithms above requires each base learner to be weak in the sense that it extracts only a small amount of structure in the density ratio function. Similar to supervised boosting, we can regularize the contribution of each learner to the estimate by introducing a learning rate denoted by $\nu\in(0,1]$. If the current estimate is $w_{k-1}$ and the optimal learner obtained in the $k$th step is $f_{k}$, we can combine them so that
$\log w_{k}=\log w_{k-1}+\nu f_{k}.$
It is known that smaller values of the learning rate $\nu$ lead to better generalization performance in both supervised learning (Friedman, 2001; Hastie et al., 2009) and unsupervised learning (Awaya and Ma, 2024). Accordingly, in our implementation, we follow them by setting $\nu$ to, for example, $0.01$.
An additional, complementary strategy for regularization is to introduce a restriction on the size of trees to avoid constructing overly complicated partition structures, so we set the maximum resolution (depth) of the trees to be, for example, between 4 and 6. Finally, the number of trees in the ensemble also needs be carefully selected. We find that choosing this number via cross-validation yields satisfactory performance.
3 Generalized Bayesian inference on additive trees for density ratios
3.1 Balancing loss as a pseudo-likelihood
To achieve uncertainty quantification on the density ratio, we use the balancing loss as a pseudo-likelihood within the framework of the generalized Bayesian inference (Bissiri et al., 2016). Specifically, we define a pseudo-likelihood using the finite-sample loss $l_{n}(w)$, that is
$L_{n,\tau}(w)=\exp(-n\tau l_{n}(w)),$
where $\tau>0$ is a “temperature” parameter. Then, given the prior distribution of $w$, $\pi_{0}(w)$, the posterior distribution of $w$ is defined as
$\pi(w)\propto\pi_{0}(w)L_{n,\tau}(w),\$
Next we describe prior specification on $w$ for efficient MCMC and a data-driven approach for setting the value of $\tau$.
3.2 Conjugate prior specification
Recall our model given in (4), we describe a Gibbs sampling strategy that iteratively samples the full conditional of each basis function $f_{k}$, given all the others. That is, how one may sample the node parameters $\beta_{k}(A)$ and the tree partition $T_{k}$ in that defines $f_{k}$ as in (5).
3.2.1 Updating leaf node parameters
To derive the full conditional for $\beta_{k}(A)$, we define the balancing weight from which $f_{k}$ is subtracted:
$\log w_{-k}:=\log w-f_{k}.$
The portion in the likelihood $L_{n,\tau}(w)$ that depends on $\beta_{k}(A)$ is given as
$L_{n,\tau}(w)\propto\exp\left(-\zeta^{-1}\tau\gamma_{k}(A)^{-1}\sum_{x^{(0)}_{i}\in A}w_{-k}^{-1}(x^{(0)}_{i})-(1-\zeta)^{-1}\tau\gamma_{k}(A)\sum_{x^{(1)}_{i}\in A}w_{-k}(x^{(1)}_{i})\right)$,
where $\zeta=n_{0}/n$ and $\gamma_{k}(A)=e^{\beta_{k}(A)}$. We note that this pseudo-likelihood resembles an exponential family kernel for $\gamma_{k}(A)$, and as such we adopt an inverse-Gaussian prior for $\gamma_{k}(A)$, which is conjugate to this kernel. Specifically, the prior $\mathrm{InverseGauss}(\mu_{k}(A),\lambda_{k}(A))$ on $\gamma_{k}(A)$ takes the following form:
$p(\gamma_{k}(A)=z\mid\mu_{k}(A),\lambda_{k}(A))$ $=\sqrt{\frac{\lambda_{k}(A)}{2\pi z^{3}}}\exp\left(-\frac{\lambda_{k}(A)(z-\mu_{k}(A))^{2}}{2\mu_{k}(A)^{2}z}\right)$
$\propto\sqrt{\frac{\lambda_{k}(A)}{2\pi z^{3}}}\exp\left(-\frac{\lambda_{k}(A)}{2\mu_{k}(A)^{2}}z-\frac{\lambda_{k}(A)}{2z}\right).$
It is straightforward to see that the conditional posterior is analytically obtained in the form of $\mathrm{InverseGauss}(\mu^{\prime}_{k}(A),\lambda^{\prime}_{k}(A))$, where
$\mu^{\prime}_{k}(A)=\sqrt{\frac{\lambda_{k}(A)+2\zeta^{-1}\tau\sum_{x^{(0)}_{i}\in A}w_{-k}^{-1}(x^{(0)}_{i})}{\frac{\lambda_{k}(A)}{\mu_{k}(A)^{2}}+2(1-\zeta)^{-1}\tau\sum_{x^{(1)}_{i}\in A}w_{-k}(x^{(1)}_{i})}},\ \lambda^{\prime}_{k}(A)=\lambda_{k}(A)+2\zeta^{-1}\tau\sum_{x^{(0)}_{i}\in A}w_{-k}^{-1}(x^{(0)}_{i})$. (6)
We set $\mu_{k}(A)$ to $1$ so that the induced marginal prior on $\log w$ can be around 0 (more on this below). The precision parameter $\lambda_{k}(A)$ induces a shrinkage toward the prior mean 1, as seen in the posterior mean $\mu^{\prime}_{k}(A)$ given in Eq (6). We set $\lambda_{k}(A)$ according to the number of trees $K$:
$\lambda_{k}(A)=\lambda_{0}K,$
where $\lambda_{0}$ is a hyperparameter. The prior value of $\lambda_{k}(A)$ is linear in $K$ because the results provided in Whitmore and Yalovsky (1978) imply that the variance of $f_{k}(x)$ for each point $x$ ($x\in\Omega$) is $O(K^{-1})$ under our model specification (see Supplementary Materials A for more detail). This choice of $\lambda_{k}(A)$ is also analogous to the amount of shrinkage recommended for the BART model (Chipman et al., 2010). We set $\lambda_{0}$ to 5 as a default value in all of our numerical experiments.
Another consideration is that the inverse Gaussian distribution has an asymmetric distribution structure, so it can also introduce skewness to the prior of the log-balancing weight $\log w$. To avoid this issue, we introduce the inverse Gaussian prior to $\gamma_{k}(A)$ and $\gamma^{-1}_{k,A}$ alternately, that is,
$\gamma_{k}(A)$ $\sim\mathrm{InverseGauss}(1,\lambda_{0}K)\ (k\text{: odd}),$
$\gamma_{k}(A)^{-1}$ $\sim\mathrm{InverseGauss}(1,\lambda_{0}K)\ (k\text{: even}).$
This also makes the marginal prior of $\log w$ exactly centered at 0 whenever the total number of trees $K$ is even.
3.2.2 Updating tree structures
We adopt the standard Bayesian CART prior (Chipman et al., 1998) on the tree $T_{k}$ as in BART (Chipman et al., 2010) and describe the prior on $T_{k}$ in terms of a recursive partition process. Starting from the full sample space as the root node, each node $A$ in the tree with depth $\mathrm{depth}(A)$ (which is 0 for the root) is further split with probability
$a_{T}(1+\mathrm{depth}(A))^{-b_{T}},\ a_{T}\in(0,1),b_{T}\in[0,\infty).$
We set the hyperparameters to the widely used default choices as $a_{T}=0.95$ and $b_{T}=2$ (Chipman et al., 2010). When splitting the node, one of the $d$ candidate dimensions of the sample space is selected with uniform probabilities, and a cut point is selected uniformly over the range in that dimension. Following the standard recipe for BART (Chipman et al., 2010; Kapelner and Bleich, 2016; Sparapani et al., 2021), we update the tree $T_{k}$ by Metropolis-Hastings moves. The details on the tree updating moves are provided in Supplementary Materials B.
3.3 Tuning the temperature $\tau$
In generalized Bayesian inference, the temperature parameter $\tau$ controls the strength of the loss-based likelihood for the balancing weight function $w$ as opposed to the prior and significantly influences the resulting posteriors, such as posterior variances and credible intervals. There is a growing number of discussions on strategies for selecting the temperature parameter empirically. Some proposed methods include information matching (Holmes and Walker, 2017), calibrating frequentist coverage rates of credible intervals (Syring and Martin, 2019), and the SafeBayes algorithm (Grünwald and van Ommen, 2017). It is not easy to adapt these methods in the current context where the task is unsupervised with an unknown function of interest specified nonparametrically.
Instead, we adopt a hierarchical Bayesian approach by placing a prior on the temperature $\tau$, following Bissiri et al. (2016); Rigon et al. (2023). The full conditional of $\tau$ is proportional to
$\pi_{0}(\tau)\tau^{n}\prod^{n_{0}}_{i=1}\exp\left(-\zeta^{-1}\tau w^{-1}(x^{(0)}_{i})\right)\prod^{n_{1}}_{i=1}\exp\left(-(1-\zeta)^{-1}\tau w(x^{(1)}_{i})\right)$,
where $\pi_{0}(\tau)$ corresponds to the prior for $\tau$. We adopt a conjugate prior $\pi_{0}(\tau)$, namely $\mathrm{Gamma}(a_{0,\tau},b_{0,\tau})$, so the full conditional of $\tau$ is $\mathrm{Gamma}(a_{1,\tau},b_{1,\tau})$, where
$a_{1,\tau}=a_{0,\tau}+n,\ b_{1,\tau}=b_{0,\tau}+nl_{n}(w).$
The analytic form of the full conditional elucidates the role $\tau$ plays and justifies the data-dependent approach to tuning of its value. Specifically, $\tau$ can be interpreted as a regularization parameter that penalizes poor overlap of the two samples. To see this, note that the full conditional mean of $\tau$ is $\frac{a_{0,\tau}+n}{b_{0,\tau}+nl_{n}(w)}$, so when the total sample size is large, the inverse of this quantity approximates the realized value of the loss $l_{n}(w)$. When the two samples have little overlap, for example, the estimated weights $w^{-1}$ for the first group and $w$ for the second group tend to be both small and close to zero, leading to a small $l_{n}(w)$ and thus a large $\tau$. Conversely, when the samples have good overlap, the weights are closer to 1, yielding a larger $l_{n}(w)$ and a smaller $\tau$. As an illustration, Figure 1 shows the posterior distributions of $\tau$ under the different levels of overlap under a simple simulation experiment. The observations ($n_{0}=n_{1}=200$) are generated from the standard normal distribution, and the second group is mean-shifted by $1,2$, and $3$, corresponding to decreasing levels of overlap. We can see that poorer overlap of the two samples leads to larger absolute values of $\log w$, along with larger values of $\tau$ corresponding to the smaller realized value of the loss $l_{n}(w)$.
Figure 1: Posterior distributions of $\tau$ under different levels of overlap. First row: The black and red points correspond to the two samples, and the dashed lines indicate the posterior means of the log-density ratios. Second row: The density functions of the posterior distributions of $\tau$ are plotted.
4 Numerical Experiments
We conduct simulation experiments to evaluate the proposed methods—both tree boosting and Bayesian additive trees. We begin with scenarios involving bivariate distributions, and then consider higher-dimensional (20-dimensional) cases. We provide a comparison with several existing approaches. The first approach to be compared with is the “density-ratio trick” by inverting a binary classifier. In particular, we adopt AdaBoost as binary classification for it uses the same class of an additive tree ensemble to classify the two classes using essentially the exponential loss (Friedman et al., 2000). The difference between this approach and our boosting algorithms will essentially be due to the different loss function we adopt. To apply the density-ratio trick, we first estimate the classification probabilities and then transform them into the density ratios by multiplying the sample size ratios. For implementation, we use the ada function provided in the R package ada (Culp et al., 2016). In addition, besides the “density-ratio trick”, we also compare our methods with two popular kernel-based density ratio estimators, namely KLEIP (Sugiyama et al., 2008) and uLSIF (Kanamori et al., 2009), implemented with the R package densratio (Makiyama, 2019).
For boosting, we assess both the FS and the GB algorithms. The learning rate $\nu$ is set to 0.01, and the maximum resolution of each tree is set to 4. The optimal number of trees is selected with the 5-fold cross-validation with the maximum number allowed 1000. For a fair comparison, AdaBoost was implemented under the same tuning parameters. For the generalized Bayesian additive tree model, we set the size of the ensemble to 200, following the standard implementation of the BART (Chipman et al., 2010). The MCMC is initialized with trees with only roots, and the size of the burn-in and the posterior sample are set to 2000 and 1000, respectively. In both types of the proposed models, we introduce 31 cut points with equal gaps in each dimension. Further details on the implementation are provided in Supplementary Materials C.
For each scenario, we evaluate the performance under two sample size settings: (i) $n_{0}=n_{1}=5000$ (“balanced” setting) and (ii) $n_{0}=9000$, $n_{1}=1000$ (“unbalanced” setting). As a measure of the accuracy, we use the following symmetrized squared error for the true density ratio $r^{*}$ and the estimated ratio $\hat{r}$:
$\frac{1}{n_{0}}\sum^{n_{0}}_{i=1}\left(\log r^{*}(x_{i}^{(0)})-\log\hat{r}(x_{i}^{(0)})\right)^{2}+\frac{1}{n_{1}}\sum^{n_{1}}_{i=1}\left(\log r^{*}(x_{i}^{(1)})-\log\hat{r}(x_{i}^{(1)})\right)^{2},$
and for each setting we calculate the MSE based on 50 simulated data sets.
4.1 2D simulations
We simulate data sets using the following three scenarios with various types of differences in the distributions of $X_{0}$ and $X_{1}$:
1. Global shift
$X_{0}\sim N\left(\mu^{(0)},I_{2}\right),\ X_{1}\sim N\left(\mu^{(1)},I_{2}\right),$
where $\mu^{(0)}=(-0.5,-0.5)^{\prime}$ and $\mu^{(1)}=(0.5,0.5)^{\prime}$.
2. Local shift
$X_{0}\sim\frac{1}{5}N(\mu_{1},\Sigma_{1})+\frac{1}{5}\sum^{5}_{k=2}N(\mu_{k},\Sigma_{k}),\ X_{1}\sim\frac{1}{5}N(\mu_{1}+\delta,\Sigma_{1})+\frac{1}{5}\sum^{5}_{k=2}N(\mu_{k},\Sigma_{k}),$
where $\delta=(0,1)^{\prime}$, and $\mu_{k}$ and $\Sigma_{k}$ ($k=1,\dots,5$) are provided in Supplementary Materials C.
3. Local dispersion difference
$X_{0}\sim\frac{1}{3}N(\mu_{1},\Sigma_{1})+\frac{1}{3}\sum^{3}_{k=2}N(\mu_{k},\Sigma_{k}),\ X_{1}\sim\frac{1}{3}N(\mu_{1},\Delta\Sigma_{1})+\frac{1}{3}\sum^{3}_{k=2}N(\mu_{k},\Sigma_{k}),$
where $\Delta=\mathrm{diag}(0.36,1)$, and $\mu_{k}$ and $\Sigma_{k}$ ($k=1,\dots,5$) are provided in Supplementary Materials C.
The simulated data sets are visualized in Figure 2. The difference between the two samples is given on a global scale in the first scenario, whereas the discrepancy is more localized in the second and third scenarios.
A comparison of the MSEs is provided in Table 1, and the result indicates that our proposed methods (the boosting and the Bayesian additive trees) overall outperform the other approaches. We note that, for the performance of the classification approach with AdaBoost, the MSEs become substantially larger when comparing the balanced and unbalanced cases in terms of the sample sizes. This confirms the common experience that inverting a binary classifier to attain the density ratio loses its efficacy when the sample sizes are unbalanced. On the other hand, the proposed balancing loss is shown to be robust even with unbalanced sample sizes.
Figure 2: A sample of simulations generated from the three scenarios created for the two-dimensional experiments. The sample sizes $n_{0}$ and $n_{1}$ are both set to 5000.
Global Shift Local Shift Local Dispersion
$n_{0}/(n_{0}+n_{1})$ 0.5 0.9 0.5 0.9 0.5 0.9
GB 0.033 0.063 0.035 0.067 0.109 0.133
(0.001) (0.002) (0.001) (0.002) (0.002) (0.005)
FS 0.036 0.071 0.035 0.070 0.111 0.133
(0.001) (0.002) (0.001) (0.002) (0.002) (0.005)
BAT 0.012 0.031 0.049 0.067 0.116 0.131
(0.000) (0.001) (0.001) (0.001) (0.002) (0.005)
DRT (AdaBoost) 0.199 2.715 0.151 3.289 0.173 2.692
(0.022) (0.250) (0.011) (0.182) (0.010) (0.199)
KLIEP 0.151 0.159 0.301 0.316 0.305 0.308
(0.006) (0.005) (0.004) (0.005) (0.004) (0.007)
uLSIF 1.348 0.470 0.412 0.390 0.130 0.199
(0.195) (0.090) (0.018) (0.026) (0.005) (0.009)
Table 1: A comparison of the MSEs and their standard errors in the 2-dim scenarios. The methods are GB (gradient boosting), FS (forward-stagewise), BAT (Bayesian additive trees), DRT (density-ratio trick based on AdaBoost), and two kernel-based methods KLIEP and uLSIF.
Figure 3 compares the estimated log-density ratios in the local shift scenario with the balanced sample size ($n_{0}=n_{1}=5000$). The estimated functions obtained with the proposed methods successfully capture the local difference structure, and importantly, the log-density ratios are significantly positive or negative for many observations included in the mixture component that has the local shift difference. For example, in the plot of the 2.5% quantiles, we find observations with yellow/red colors, which implies the log-density ratios are significantly larger than 0, mostly in the corresponding mixture component. Similar results are obtained for the other scenarios and sample size settings, and we provide their visualizations in Supplementary Materials D.
Figure 3: The estimated log-density ratios obtained with the algorithms considered in the two-dimensional examples for the local shift scenario ($n_{0}=n_{1}=5000$). For the Bayesian additive model, the 2.5% quantiles and 97.5% quantiles of the log-density ratios evaluated point-wise are also displayed. The difference between FS and GB is so small that we present only the results for GB.
4.2 Higher-dimensional settings
In this experiment, we evaluate the performance of the density ratio estimation algorithms in higher-dimensional settings. We generate 20-dimensional observations from essentially a latent factor model. First, we simulate latent variables $Z$ from 4-dimensional distributions, in which the difference between the two samples is introduced, and then project them into the 20-dimensional space as
$\Lambda Z+\epsilon_{i},\ \epsilon_{i}\sim N(\mathbf{0},(0.1)^{2}I_{20}),$
where $\Lambda$ is a randomly generated orthonormal $20\times 4$ matrix satisfying $\Lambda^{\prime}\Lambda=I_{4}$. The distributions that the latent variables follow for the two groups, indicated by random variables $Z_{0}$ and $Z_{1}$, are defined under the two scenarios described below.
1. Location Shift
$Z_{0}\sim\frac{4}{5}N(\mu_{1},\Sigma_{1})+\frac{1}{5}N(\mu_{2},\Sigma_{2}),\ Z_{1}\sim\frac{4}{5}N(\mu_{1}+\delta,\Sigma_{1})+\frac{1}{5}N(\mu_{2},\Sigma_{2}),$
where $\delta=(1,0,0,0)^{T}$. The parameters $\mu_{k}$ and $\Sigma_{k}$ ($k=1,2$) are provided in Supplementary Materials C.
2. Dispersion Difference
$Z_{0}\sim\frac{4}{5}N(\mu_{1},\Sigma_{1})+\frac{1}{5}N(\mu_{2},\Sigma_{2}),\ Z_{1}\sim\frac{4}{5}N(\mu_{1},\Delta\Sigma_{1})+\frac{1}{5}N(\mu_{2},\Sigma_{2}),$
where $\Delta=\mathrm{diag}(0.5,1,1,1)$. The parameters $\mu_{k}$ and $\Sigma_{k}$ ($k=1,2$) are given in Supplementary Materials C.
The orthonormal matrix $\Lambda$ is generated with the randortho function in the R package pracma (Borchers, 2023).
Figure 4 shows the distribution of the first two latent factors, which account for the differences between the two groups, along with the corresponding true and estimated log density ratios. The latent factor model for the true densities presents a challenging case for density ratio estimation, and in particular for tree-based methods that use axis-aligned partitions along the original dimensions of the sample space. Nevertheless, as shown in Table 2, the proposed methods still produce more accurate estimates, and the performance is robust to the unbalanced sample sizes. But even more importantly, in such challenging settings, uncertainty quantification becomes even more important to draw robust conclusions. The inferred density ratios must be gauged against appropriate measures of uncertainty to reach reliable inference. Our generalized Bayesian approach serves this need—the yellow/red regions in the lower quantiles and blue points in the upper quantiles are those that fall outside the 95% credible bands, and they indeed correspond to the regions with true differences. Had this uncertainty quantification been unavailable, the inferred different regions would have included much larger portions of the sample space under all of the estimates.
Location Shift Dispersion
$n_{0}/(n_{0}+n_{1})$ 0.5 0.9 0.5 0.9
GB 0.073 0.125 0.150 0.228
(0.001) (0.002) (0.003) (0.005)
FS 0.076 0.136 0.158 0.237
(0.001) (0.002) (0.004) (0.006)
BAT 0.059 0.115 0.171 0.202
(0.001) (0.002) (0.004) (0.003)
DRT (AdaBoost) 0.237 3.858 0.151 3.031
(0.027) (0.119) (0.005) (0.157)
KLIEP 0.611 0.609 0.547 0.548
(0.006) (0.008) (0.001) (0.001)
uLSIF 3.379 3.621 0.546 0.547
(0.074) (0.089) (0.001) (0.001)
Table 2: A comparison of the MSEs and their standard errors in the 20-dim scenarios. The methods are GB (gradient boosting), FS (forward-stagewise), BAT (Bayesian additive trees), DRT (density-ratio trick based on AdaBoost), and two kernel-based methods KLIEP and uLSIF.
Figure 4: A comparison of the true log-density ratios and the estimated ratios projected to the space of the latent variables $(z_{1},z_{2})$, where the result for the location shift scenario is shown ($n_{0}=n_{1}=5000$). The difference in the results for FS and GB are so small that we only present those for GB.
5 Assessing the quality of synthesized microbiome compositions
In recent years, various generative models have been introduced to imitate the sampling mechanisms of various complex biomedical data. In particular, several approaches have been proposed for generating microbiome compositions (Sankaran et al., 2025). A common need is to assess the quality of the synthetic samples drawn from these models in comparison with the actual observed data. We apply our two-sample comparison framework to synthetic samples simulated from different generative models and the corresponding observed samples, thereby assessing the data-generation quality of these models.
The real microbiome data are obtained from the open-source Curated Metagenomic Data (Pasolli et al., 2017), available as an R package curatedMetagenomicData. We focus on the relative abundance data and apply the following subsetting rules to the relative abundance data: removing samples whose microbial relative abundance does not sum up to 1 (i.e., $\geq 0.99$); then, keep taxa with average relative abundance $\geq 0.1\%$; among those taxa, keep only samples that have zeros in fewer than 10% of taxa and whose mean abundance is at least the overall average. After subsetting, each row contains 123 columns of species-level relative abundance. We use one of the largest studies, the IBDMDB dataset (Lloyd-Price et al., 2019) in the database, as the data of interest. (The samples are in fact collected at multiple time points for each given subject, but the longitudinal nature of the data is important for scientific analysis but could be ignored for the purpose of this demonstration since we will only be focusing on assessing how well generative models can approximate the marginal distribution of the samples.) We split the data into training ($n=1166$) and testing ($n=292$) sets. The generative models are trained using the training data, and the testing data are used in the follow-up two-sample comparison with the synthetic samples. We include various popular generative models, including both classical parametric models and state-of-the-art deep neural network-based models. The parametric models considered are the Dirichlet model (Holmes et al., 2012; La Rosa et al., 2012) and the Dirichlet tree model (Dennis III, 1991; Wang and Zhao, 2017; Mao and Ma, 2022). The nonparametric generators based on neural networks include continuous normalizing flows through conditional flow matching (Lipman et al., 2022; Tong et al., 2023, ICFM) (code available on GitHub), and a generative adversarial network model called MB-GAN (Rong et al., 2021) (available on Github), which is tailored for microbiome data generation. Our objective here is not to argue which model is better than others, as such a judgment needs to be context-specific and requires more careful tuning of each model. We aim to simply provide an example that density ratio-based two-sample comparison can provide insights into how various generative models can be gauged given the observations that they generate.
Train v.s. Generated
Test v.s. Generated
Figure 5: Principal Coordinates Analysis (PCoA) plots based on Bray-Curtis distance between samples. The blue and orange points represent the real sample and the generated sample, respectively. First row presents the training sample vs the generated sample; second row presents the test sample vs the generated sample.
The PCoA plots based on the Bray-Curtis distance (Bray and Curtis, 1957) in Figure 5 provide a low-dimensional visualization to compare the generated sample and the train/test sample. These plots clearly reveal that the parametric generators (Dirichlet and Dirichlet-Tree) have a smaller coverage in the support than the train/test sample, while the two neural network-based approaches are better in this regard. However, it is hard to further differentiate the quality of the two neural network-based generators based on such visualization.
We apply the generalized Bayesian additive trees and obtain a posterior sample of the density ratio evaluated at each of the real and the synthetic samples. We then computed the posterior mean along with a pointwise 95% credible band for the density ratio, which are presented in Figure 6 (training sample vs synthetic sample) and Figure 7 (testing sample vs synthetic sample). The color reflects the estimated (log) density ratio. The estimated density ratios in both Figures 6 and 7 provide a clear contrast of the generation quality compared to the true sample. The estimated log-density ratio for the MB-GAN-generated sample versus the real sample is close to 0 across the support. The credible bands provided by our proposed method offer uncertainty quantification for the density ratio, providing further evidence that 95% of the posterior mass of the log density ratio comfortably covers 0 over most of the sample space, which is not the case for the parametric models and ICFM.
Train v.s. Generated: Posterior mean of DRE
Train v.s. Generated: Lower 2.5% MCMC quantile of DRE
Train v.s. Generated: Upper 97.5% MCMC quantile of DRE
Figure 6: PCoA plots colored by the density ratio estimates of the training sample density over the generated sample density, and the 2.5% and 97.5% credible bands.
Test v.s. Generated: Posterior mean of DRE
Test v.s. Generated: Lower 2.5% MCMC quantile of DRE
Test v.s. Generated: Upper 97.5% MCMC quantile of DRE
Figure 7: PCoA plots colored by the density ratio estimates of the testing sample density over the generated sample density, and the 2.5% and 97.5% credible bands.
6 Conclusion
We have introduced a new approach to learning the density ratio between two i.i.d. samples using an additive tree model as the function class. We showed that a new loss function called balancing loss can be very useful in learning the density ratio based on this function class. Not only does it provide effective boosting algorithms to attain functional (point) estimates, but it also provides a convenient means to generalized Bayesian inference on the density ratio, attaining uncertainty quantification which can be critical in challenging problems where the uncertainty in the learned density ratio can be substantial. The training algorithms we presented, both for boosting and for generalized Bayesian learning, are only minor modifications of their counterparts widely used in supervised learning based on this model class, and so their implementation would involve only minor modifications of existing software for boosting and BART models.
Additive tree models do have their limitations. In particular, they use axis-aligned partitions and therefore can sometimes be less effective in capturing high-order interactions, and they generally do not perform well on very high-dimensional data, i.e., those involving more than hundreds of dimensions. In such scenarios, an effective strategy is to couple the tree-based approaches with strategies for learning effective representations of the data. Examples of such representations might include low-dimensional subspace assumptions (e.g., manifolds) or other highly non-linear embeddings through state-of-the-art neural network-based encoders. Even in such cases, we believe that learning the density ratio between two samples is a worthy objective and will typically be easier to accomplish than learning the density of each of the two samples per se.
7 Acknowledgment
This research is funded in part by NIGMS grant R01-GM135440 and NSF grant DMS-2152999. Part of the research was completed when LM and YX were at Duke University. NA’s research is supported by the Japan Society for the Promotion of Science KAKENHI (grant numbers: 25K16618, 24H00142, and 25H00546). The authors thank Dr. Ganchao Wei for preprocessing the IBDMDB data and generating the synthetic samples.
Supplementary Materials
Appendix A Proofs
A.1 Proof of Proposition 2.1
We first fix the finite tree partition $T_{k}$. After adding the piece-wise constant function $f_{k}$, the loss is given as
$l(w)=\sum_{A\in T_{k}}\left[e^{-\beta_{k}(A)}\int w^{-1}_{k-1}pd\mu+e^{\beta_{k}(A)}\int w_{k-1}qd\mu\right]$
By the relationship between the arithmetic mean and the geometric mean, each summand is bounded below by
$2\sqrt{\int w^{-1}_{k-1}pd\mu\int w_{k-1}qd\mu},$
and the equation is attained if
$e^{-\beta_{k}(A)}\int w^{-1}_{k-1}pd\mu=e^{\beta_{k}(A)}\int w_{k-1}qd\mu,$
and this condition leads to the first result on the optimal $\beta_{k}(A)$. Also, when the lower bound is achieved for all $A\in T_{k}$, the loss can be written as
$l(w)=2\sqrt{\int w^{-1}_{k-1}pd\mu\int w_{k-1}qd\mu},$
and minimizing it with respect to $T_{k}$ is equivalent to maximizing the discrete Hellinger loss given in the claim. ∎
Also, we can derive the following corollary, by setting $T_{k}$ to the root tree.
Corollary A.1.
Let $w$ be the current estimate of the balancing function and $c$ be the constant defined as $c=\sqrt{\mathbb{E}_{p}[w^{-1}]/\mathbb{E}_{q}[w]}$ and thus satisfies
$\mathbb{E}_{p}[(cw)^{-1}]=\mathbb{E}_{q}[(cw)].$
Then we obtain an improvement in the balancing loss as
$l(w)\geq l(cw).$
A.2 Derivation of the regularized prior for the node parameters
According to Whitmore and Yalovsky (1978), when $W\sim\mathrm{InverseGauss}(1,\theta)$ and $\theta$ are large, the distribution of the following variable is approximated by the standard normal distribution:
$\frac{1}{2\sqrt{\theta}}+\sqrt{\theta}\log W.$
Hence, when $\gamma_{k}(A)=e^{\beta_{k}(A)}$ follows $\mathrm{InverseGauss}(1,K\lambda_{0})$, the leaf node parameter $\beta_{k}(A)$ approximately follows
$N\left(-\frac{1}{2\sqrt{K\lambda_{0}}},\frac{1}{K\lambda_{0}}\right).$
This observation implies that, under the prior such that
$\gamma_{k}(A)\sim\mathrm{InverseGauss}(1,K\lambda_{0})\ (k:\text{odd}),\ \gamma_{k}(A)^{-1}\sim\mathrm{InverseGauss}(1,K\lambda_{0})\ (k:\text{even}),$
the distribution of $\log w$ is approximated by $N(0,\lambda_{0}^{-1})$.
Appendix B Details of the posterior sampling for the Bayesian additive tree model
B.1 MCMC algorithm
The Bayesian additive model include three kinds of unknown variables: (i) tree partition structures $T_{1},\dots,T_{K}$, (ii) leaf node parameters $\beta_{1},\dots,\beta_{K}$, where $\beta_{k}$ is the collection of the parameters defined for the $k$th basis function, and (iii) the temperature $\tau$. Their posterior distributions are estimated with the following Metropolis-within-Gibbs algorithm:
  1. 1.
    For $k=1,\dots,K$,
    1. (a)
      Compute the residual as $\log w_{-k}=\sum_{l\neq k}f_{l}$.
    2. (b)
      Update $T_{k}\mid w_{-k},\tau$.
    3. (c)
      Update $\beta_{k}\mid w_{-k},T_{k},\tau$.
  2. 2.
    Update $\tau\mid w$, where $w$ is the current balancing weight.
Updating the node parameters $\beta_{k}$ and the temperature $\tau$ is straightforward due to the conjugate structures. The algorithm to update $T_{k}$ with Metropolis-Hasting is detailed in the next section.
B.2 Updating trees
In this section, we consider updating the $k$th tree denoted by $T_{k}$ given the other parameter values with the proposals originally adopted for the Bayesian CART model (Chipman et al., 1998), and the details are obtained by following Kapelner and Bleich (2016).
In each iteration of the MCMC, we propose a new tree structure $T^{*}_{K}$ by altering the current tree $T_{k}$ with one of the three moves, GROW, PRUNE, and CHANGE, defined as follows:
GROW:
Randomly select one leaf node and split it to generate two child nodes.
PRUNE:
Randomly select one second-generation internal (2GI) node, that is, a node with two child nodes that are terminal, and remove these child nodes.
CHANGE:
Randomly select one 2GI node and change the splitting rules by sampling one rule from the prior distribution.
One of the moves is selected stochastically according to the fixed probabilities $p(\mathrm{GROW})$, $p(\mathrm{PRUNE})$, and $p(\mathrm{CHANGE})$, and they are all set to 1/3 in our numerical experiments.
Let $T^{*}_{k}$ denote the new proposed tree structure. Then, the proposal is accepted with the probability $\min\{\alpha,1\}$, where
$\alpha=\underbrace{\frac{p(T^{*}_{k})}{p(T_{k})}}_{\text{Prior}}\underbrace{\frac{p(T^{*}_{k}\to T_{k})}{p(T_{k}\to T^{*}_{k})}}_{\text{Transition}}\underbrace{\frac{p(\mathrm{data}\mid T_{k},p(\{w_{-k}(x_{i})\}^{n}_{i=1})}{p(\mathrm{data}\mid T^{*}_{k},p(\{w_{-k}(x_{i})\}^{n}_{i=1})}}_{\text{Loss-based likelihood}}$
and $\{w_{-k}(x_{i})\}^{n}_{i=1}$ are the residuals. If a node $A$ is selected for updating $T_{k}$, then the ratios corresponding to the priors and the transition probabilities for the three moves are given as
[…]
(see Kapelner and Bleich (2016) for the details.)
For calculating the likelihood ratios, we need to note that we can analytically evaluate the likelihood with the leaf node parameters integrated out due to the conjugate structure as follows:
$\int p(\beta_{k})L_{n,\tau}(w)d\beta_{k}\propto\prod_{A\in\mathcal{L}(T_{k})}L_{k,A},$
where $p(\beta_{k})$ is the joint prior density of the leaf node parameters, $\mathcal{L}(T_{k})$ is the collection of the leaf nodes, and
$L_{A}=\sqrt{\frac{\lambda_{k}(A)}{\lambda^{\prime}_{k}(A)}}\exp\left(\frac{\lambda_{k}(A)}{\mu_{k}(A)}-\frac{\lambda^{\prime}_{k}(A)}{\mu^{\prime}_{k}(A)}\right).$
Hence, for the GROW move in which the node $A$ is split into $A_{l}$ and $A_{r}$, the likelihood ratio is given as
$\frac{L_{A_{l}}L_{A_{r}}}{L_{A}},$
while in the PRUNE move that removes $A_{l}$ and $A_{r}$, the ratio is the inverse. The CHANGE move generates two new children $A^{*}_{l}$ and $A^{*}_{r}$, so the likelihood ratio is given as
$\frac{L_{A^{*}_{l}}L_{A^{*}_{r}}}{L_{A_{l}}L_{A_{r}}}.$
Appendix C Details of experiments
C.1 2D numerical experiments
The parameters for simulating data sets are specified as follows:
2. Local shift
$\mu_{1}=(9.0,9.9)^{\prime}$, $\mu_{2}=(-2.5,1.4)^{\prime}$, $\mu_{3}=(-2.3,-9.7)^{\prime}$, $\mu_{4}=(3.4,5.9)^{\prime}$, $\mu_{5}=(5.8,-9.5)^{\prime}$, $(\Sigma_{1}(1,1),\Sigma_{1}(1,2),\Sigma_{1}(2,2))=(2.9,0.5,1.1)$, $(\Sigma_{2}(1,1),\Sigma_{2}(1,2),\Sigma_{2}(2,2))=(1.2,-0.6,2.8)$,
$(\Sigma_{3}(1,1),\Sigma_{3}(1,2),\Sigma_{3}(2,2))=(2.3,-1.0,1.7)$, $(\Sigma_{4}(1,1),\Sigma_{4}(1,2),\Sigma_{4}(2,2))=(1.1,-0.4,2.9)$,
and $(\Sigma_{5}(1,1),\Sigma_{5}(1,2),\Sigma_{5}(2,2))=(3.0,0.2,1.0)$.
3. Local dispersion difference
$\mu_{1}=(1.9,-7.2)^{\prime}$, $\mu_{2}=(-5.7,5.3)^{\prime}$, $\mu_{3}=(-2.3,-1.5)^{\prime}$, $\mu_{4}=(7.5,-3.1)^{\prime}$, $\mu_{5}=(3.1,9.5)^{\prime}$, $(\Sigma_{1}(1,1),\Sigma_{1}(1,2),\Sigma_{1}(2,2))=(1.0,-0.4,0.8)$,
$(\Sigma_{2}(1,1),\Sigma_{2}(1,2),\Sigma_{2}(2,2))=(1.3,0.7,2.7)$, $(\Sigma_{3}(1,1),\Sigma_{3}(1,2),\Sigma_{3}(2,2))=(1.0,0,3.0)$,
$(\Sigma_{4}(1,1),\Sigma_{4}(1,2),\Sigma_{4}(2,2))=(2.9,0,1.1)$, and $(\Sigma_{5}(1,1),\Sigma_{5}(1,2),\Sigma_{5}(2,2))=(2.4,-0.9,1.6)$.
C.2 20D dimensional experiments
$\mu_{1}$ is set to $(-0.5,0,0,0)^{\prime}$ in the location shift scenario and $(0,0,0,0)^{\prime}$ in the dispersion difference scenario. For the other parameters, $\mu_{2}=(0,0,0,0)^{\prime}$, $\Sigma_{1}=I_{4}$, and $\Sigma_{2}=4I_{4}$ in both scenarios.
C.3 Notes on the implementation of the density ratio estimators
In the proposed boosting algorithms, we prune leaf nodes that include only observations from one group to avoid division by zero in computing the optimal values of leaf node parameters.
For AdaBoost, we used the ada function from the ada package (Culp et al., 2016), and the optimal number of trees is selected via the $5$-fold cross validation based on the classification error reported by the function. We found that the selected number tends to be close to zero when the two-sample difference is small (such as the local shift scenario in the 2D experiments), so we set the minimum number of trees to 10.
For the KLIEP and uLSIF, we use the functions KLIEP and uLSIF from the R package densratio (Makiyama, 2019) with the default parameter values. It should be noted that we detected an error in the source code for the KLIEP function 111 The estimated density ratios are not correctly computed on observations in a test data set in the original package. , so we used a fixed version that is available at https://github.com/nawaya040/densratio.
Appendix D Additional figures
We provide additional visualizations of the results for the settings described in Section 4.1. Figure 8 shows the estimated log-density ratios for the first (global shift) and third (local dispersion) scenarios under the balanced sample size setting. The estimated ratios for the second (local shift) scenario under the unbalanced sample size setting are displayed in Figure 9.
Figure 8: The estimated log-density ratios obtained with the algorithms considered in the two-dimensional examples for the global shift scenario and the local dispersion scenario ($n_{0}=n_{1}=5000$). For the Bayesian additive model, the 2.5% quantiles and 97.5% quantiles of the log-density ratios evaluated point-wise are also displayed. The difference between FS and GB is so small that we present only the results for GB.
Figure 9: The estimated log-density ratios obtained with the algorithms considered in the two-dimensional examples for the global shift scenario and the local dispersion scenario ($n_{0}=9000$ and $n_{1}=1000$). For the Bayesian additive model, the 2.5% quantiles and 97.5% quantiles of the log-density ratios evaluated point-wise are also displayed. The difference between FS and GB is so small that we present only the results for GB.