### v4

# Two-sample Comparison through Additive Tree Models for Density Ratios

Abstract
The ratio of two densities provides a direct characterization of their differences. We consider the two-sample comparison problem by estimating this ratio given i.i.d. observations from two distributions. To this end, we propose additive tree models for density ratio estimation along with efficient algorithms using a new loss function, the balancing loss. The loss allows tree-based models to be trained using several algorithms originally designed for supervised learning, such as forward-stagewise optimization and gradient boosting. Moreover, the balancing loss resembles an exponential family kernel, and it can serve as a pseudo-likelihood with conjugate priors. This property enables generalized Bayesian inference on the density ratio using backfitting samplers designed for Bayesian additive regression trees (BART). Our Bayesian strategy provides uncertainty quantification for the inferred density ratio, which is critical for applications involving high-dimensional and data-limited distributions with potentially substantial uncertainty. We further show connections of the balancing loss to the exponential loss in binary classification and to the variational form of $f$-divergence, particularly the squared Hellinger distance. Numerical experiments demonstrate that our method achieves both accuracy and computational efficiency, while uniquely providing uncertainty quantification. Finally, we demonstrate its application to assessing the quality of generative models for microbiome compositional data.
Keywords: Generalized Bayesian inference, density ratio estimation, tree boosting, generative models
1 Introduction
A vast range of applications involve comparing two groups of observations to assess their differences. The classical formulation of the two-sample comparison often involves parametric or semiparametric assumptions on the underlying distributions and focuses on testing the null hypothesis of no difference with or without a given alternative in mind. However, this hypothesis testing perspective is often inadequate for most modern applications of two-sample comparison. Practitioners today typically also desire to learn the specific nature of the underlying difference. This problem lies at the core of numerous other modern statistical analyses and machine learning objectives.
For example, in large-scale biomedical studies involving the comparison of genetic measurements and various biomarkers between patients and controls, there is often no question that some types of two-sample differences do exist, and with a sufficiently large sample size, the null of no difference at all will be rejected. It is more helpful to specifically pinpoint what such differences are in order to help understand their biological implication. Another timely application is the assessment of generative models. The key to assessing the quality of different models and algorithms for generating synthetic samples that imitate the real sampling distribution of the training data lies in deciphering the distributional difference between the real sampling distribution and that induced by the generative models, which is essentially a two-sample comparison problem. Some generative models, such as generative adversarial networks (GANs) (Goodfellow et al., 2014), essentially iteratively apply two-sample comparison and use the estimated misfit between the two distributions to sequentially improve the generative model. Two-sample comparison is also a quintessential problem in causal inference, in which consistent estimation of treatment effects often relies on correctly adjusting for the difference in the covariate distributions between the treated and the control groups.
Our focus here is on carrying out a two-sample comparison in a nonparametric fashion through estimation rather than hypothesis testing, using the density ratio function as the functional “statistic” of interest for characterizing two-sample differences. The density ratio fully summarizes how two distributions differ across the sample space, and thus allows one not only to carry out hypothesis testing but also to identify the nature of the difference as well. If, for example, the estimated density ratio deviates “significantly” away from 1 on parts of the sample space, it allows us to draw statistical conclusions regarding whether, where, and how the two distributions differ. At the same time, the notion of “significant” deviation will require some form of uncertainty quantification for the inferred density ratio, which will become available under our formulation.
One may reasonably raise the concern that nonparametric density ratio learning is simply “too hard” for complex, high-dimensional data because density estimation itself in such settings is already a challenging problem. Why then would one attempt an even harder problem of learning the ratio of two densities? Though it might appear counterintuitive, we argue that density ratio estimation (DRE) is in fact a much easier problem than density estimation in most modern applications. A density by definition is always a density ratio, with the denominator being the corresponding base measure against which the density is computed. The typical base measure, for example, the Lebesgue measure or the uniform density in continuous cases, is generally chosen out of mathematical convenience and lack of prior knowledge, without taking any account of the dataset at hand (other than that the observations are continuous). This base measure generally does not resemble any sampling distribution in high-dimensional or otherwise complex problems. When good reliable knowledge of the sampling distribution does exist, one would be in the much simpler classical setting of a “goodness-of-fit” problem. In contrast, in most applications of two-sample comparison, the two distributions are often quite similar to each other and can serve as a “baseline of reference” for each other. As such, the density ratio between these two samples then is often much simpler than the underlying densities (i.e., their ratios with respect to the Lebesgue measure).
We believe this explains a well-known observation in the DRE literature, that though estimating a density ratio can be done by first separately estimating the two densities and then computing their ratios based on the density estimates, it is usually more effective to adopt a method directly aimed at the density ratio, especially when the difference exists only in a small subset of the sample space or in a subset of the dimensions. Several authors have adopted this perspective and proposed methods directly targeting the density ratio. Some approaches include kernel-based methods (e.g. Sugiyama et al., 2007b; Kanamori et al., 2009; Hido et al., 2011; Liu et al., 2013; Yamada et al., 2013) and more recently neural network-based methods (Kato and Teshima, 2021).
Beyond the stand-alone two-sample comparison problem which is our focus herein, DRE has also been applied as a component in various learning methods in which two-sample comparison arises as a sub-problem directly or indirectly. A far-from-exhaustive list includes covariate shift correction (Shimodaira, 2000; Sugiyama et al., 2007a; Reddi et al., 2015; Stojanov et al., 2019), outlier detection (Hido et al., 2011; Kato and Teshima, 2021), causal inference (Uehara et al., 2020), and change point detection (Kawahara and Sugiyama, 2009; Liu et al., 2013; du Plessis et al., 2015). Additionally, estimating density ratios has been utilized in Markov Chain Monte Carlo (MCMC) for intractable likelihoods (Kaji and Ročková, 2023).
We list several desiderata for learning the density ratio and discuss how we aim to achieve them. First, the model or function class representing the density ratio should be flexible enough to effectively approximate complex forms of the underlying truth. At the same time, it should scale well with both sample size and dimensionality. To these ends, we propose a class of density ratio learners based on additive tree ensembles (see e.g., Friedman et al., 2000; Hastie et al., 2009). Additive tree ensembles are a natural candidate for the density ratio problem, because of their success in classification problems. In fact, in the machine learning literature, a widely used technique to learn the density ratio between two samples is the so-called “density-ratio trick” (Sugiyama et al., 2012; Durkan et al., 2020; Thomas et al., 2022), which involves first fitting a binary classifier to the two samples, and then, based on the estimated posterior odds of the sample label, computing the density ratio. Additive tree boosting such as AdaBoost (Friedman et al., 2000) is among the most effective methods for binary classification.
In contrast to this popular perspective, we hold that DRE should not and does not need to involve the indirect route of inverting a binary classifier based on additive tree ensembles, but can instead be directly learned from the latter. To establish the direct route, we propose a new loss function called “balancing loss”, which has a functional form similar to the exponential loss used in the classification AdaBoost (Friedman et al., 2000) but is minimized by the true density function. The resulting estimates are often substantially more accurate than what the density-ratio trick produces, especially when the sample size is unbalanced, that is, when the setting is more challenging for classifiers that aim to maximize the total successful classification rate. We demonstrate the limitation of adopting the classification AdaBoost in the experiments.
In terms of computational strategies, we introduce two boosting algorithms for estimating the additive tree model on density ratios. We use an additive tree ensemble as the function class for approximating the density ratio, by adding tree-based weak learners. The first algorithm is based on forward-stagewise fitting (Hastie et al., 2009) and it operationally resembles AdaBoost by minimizing the balancing loss by repeatedly fitting a single-tree weak learner to reweighted observations. The second algorithm follows the general paradigm of gradient boosting (Friedman, 2001).
These algorithms inherit important advantages of supervised tree boosting — they can be implemented at a small computation cost, and at the same time achieve good numerical performance, especially when the difference lies in a marginal distribution of a subset of the variables. One can incorporate regularization to avoid overfitting effectively by making the contribution of each base learner weak through a small learning rate, following the standard approach in supervised boosting (Hastie et al., 2009; Friedman, 2001).
The second desideratum is to attain uncertainty quantification for the estimated density ratio, thereby allowing statistical inference to proceed. This aspect has mostly been ignored in the existing literature on DRE. One exception is Bayesian nonparametric models for the density ratio using a single tree model (Ma and Wong, 2011; Soriano and Ma, 2017). While they provide a means to examine posterior uncertainty in a fully Bayesian manner, piecewise-constant functions represented by single trees of controlled complexity are very restrictive and therefore inadequate for approximating complex density ratios. Also, more importantly, they introduce generative models for the two samples, and this approach can be inefficient in terms of the two-sample comparison, since estimating the common distributional structure is redundant. As such, generalizing these existing Bayesian models to build an additive tree model with which we can focus on the two-sample difference is quite challenging. To overcome this, we propose a generalized Bayesian approach (Bissiri et al., 2016) using the balancing loss to construct a pseudo-likelihood for this model class. Interestingly, we show that, due to the resemblance of the balancing loss to an exponential family kernel, there exists a natural conjugate prior for the model parameters under the pseudo-likelihood, which allows standard sampling strategies for BART models (Chipman et al., 2010) to be directly adopted for sampling the Gibbs posterior on the density ratio.
The remainder of the paper is organized as follows. Section 2 formulates the balancing loss and introduces the boosting algorithms. Section 3 proposes the generalized Bayesian inference based on the balancing loss and the BART model designed for density ratio estimation. Sections 4 and 5 present numerical experiments and an application to the microbiome data, and Section 6 concludes.
2 Boosting Algorithms for Density Ratio Estimation
In this section, we propose new boosting algorithms for density ratio estimation with a novel loss function called balancing loss. We first formulate the balancing loss and discuss its connection to other learning problems. Then, we introduce optimization procedures minimizing the loss, using the boosting algorithms (Friedman, 2001).
2.1 The Balancing Loss for Density Ratio Learning
Let $P$ and $Q$ be unknown probability measures defined on the common sample space denoted by $\Omega$. We focus on the case where $\Omega$ is a $d$-dimensional Euclidean space. We assume that the two measures $P$ and $Q$ admit density functions with respect to the Lebesgue measure $\mu$ and they are denoted by $p$ and $q$, respectively. Our goal is to estimate the unknown density ratio $p/q$ based on two i.i.d. samples.
Let $r$ denote an estimate of the density ratio. In the following discussion, we assume that $r$ is positive $\mu$-almost everywhere. To introduce an algorithm for optimizing $r$, we define a loss function called the balancing loss based on the following fact: If $r^{*}$ is equal to the true density ratio $p/q$, then for any measurable subset $B\subset\Omega$, we have
$\int_{B}(r^{*})^{-1/2}p\,d\mu=\int_{B}(r^{*})^{1/2}q\,d\mu=\int_{B}\sqrt{pq}\,d\mu.$ (1)
When $B=\Omega$, the whole sample space, this corresponds to the affinity between $p$ and $q$. The first equation implies that if we weight the two samples respectively with $(r^{*})^{-1/2}$ and $(r^{*})^{1/2}$, their weighted measures over any measurable set $B$ becomes equal. In this sense, we say that the two samples are “balanced”. Motivated by this observation, we introduce the notation $w=r^{1/2}$ to denote the balancing function, which is intended to approximate the square root of the density ratio, $\sqrt{p/q}$. We then consider the following balancing loss function:
$l(w)=\mathbb{E}_{p}[w^{-1}]+\mathbb{E}_{q}[w].$
Note that by the arithmetic-geometric mean inequality, the loss is bounded below by $2\int\sqrt{pq}\,d\mu$, and the equation holds if $w^{-1}p=wq$, or equivalently, $w=\sqrt{p/q}$. This simple justification is an advantage of selecting this functional form compared to other choices that can follow Eq. (1), such as $(\mathrm{E}_{p}[w^{-1}]-\mathrm{E}_{q}[w])^{2}$.
In practice, suppose we observe i.i.d. draws from $p$ and $q$:
$x^{(0)}_{1},\dots,x^{(0)}_{n_{0}}\sim p,\ x^{(1)}_{1},\dots,x^{(1)}_{n_{1}}\sim q.$
Then the finite-sample version of the balancing loss is defined by
$l_{n}(w)=\frac{1}{n_{0}}\sum^{n_{0}}_{i=1}w^{-1}(x^{(0)}_{i})+\frac{1}{n_{1}}\sum^{n_{1}}_{i=1}w(x^{(1)}_{i})$,
where $n=n_{0}+n_{1}$ indicates the total number of observations.
2.2 Connections to Existing Concepts in Statistical Learning
While the design of the balancing loss may appear ad hoc at first, we next show that it can be naturally motivated from two well-established perspectives in machine learning.
Connection 1: The “Density-ratio Trick” Based on Binary Classification
A first connection is to the “density-ratio trick”, which estimates the density ratio by inverting a binary classifier trained to distinguish the two samples (Sugiyama et al., 2012; Durkan et al., 2020; Thomas et al., 2022). It involves inverting a binary classifier to compute the density ratio. As we discuss below, the balancing loss arises naturally as a variant of the exponential loss, which is used for classification with AdaBoost (Friedman et al., 2000), but only the balancing loss targets the density ratio directly and thus is more robust than the density ratio trick to scenarios in which the two sample sizes are unbalanced and/or the two-sample differences exist locally in a small subset of the sample space.
To see the connection between the exponential loss and the proposed balancing loss, we suppose that, in generating the two samples, one first randomly decides which sample to draw the observation from with equal probability. That is, one first draws a random variable $y$ that takes 1 and -1 with a 50-50 chance. Given $y$ one generates the corresponding sample $x$ from $P$ if $y=1$ and from $Q$ if y=1y=-1, respectively. Then it is straightforward to show that
$\mathbb{E}_{p}[w^{-1}]+\mathbb{E}_{q}[w]\propto\mathbb{E}[\exp(-yF)],$
where $F=\log w$, and the expectation in the latter is taken with respect to the joint distribution of $x$ and $y$. What we have on the right-hand side is exactly the exponential loss (Friedman et al., 2000; Hastie et al., 2009).
We note that the exponential loss used in AdaBoost in fact integrates $y$ with respect to its marginal distribution from the data rather than using a 50-50 split. Because the population minimizer for AdaBoost is the log posterior odds between the two samples (Friedman et al., 2000), theoretically we can estimate the density ratio via the density ratio trick (Sugiyama et al., 2012), namely,
$\frac{p(x)}{q(x)}=\frac{\mathrm{Pr}(y=-1)}{\mathrm{Pr}(y=1)}\frac{\mathrm{Pr}(y=1\mid x)}{\mathrm{Pr}(y=-1\mid x)},$
where the prior ratio is replaced with the sample size ratio in practice. This approach, however, can introduce difficulties if the sample size is unbalanced. Under such imbalance, AdaBoost tends to show poor performance for classifying the observations with a small sample size, because the influence of the rare group under the empirical exponential loss $\sum_{i}e^{-y_{i}F(x_{i})}$ ($x_{i}$ and $y_{i}$ are an observation and its class, respectively) is substantially weak, which can result in a severe bias (see e.g., Sun et al., 2007, for a more in-depth discussion on this issue). Also, as a problem specific to density ratio estimation, the classification algorithm does not necessarily incorporate the information of prior odds explicitly, and multiplying the prior odds, that is, subtracting the information of the prior, can introduce a bias. An additional experiment for confirming the existence of the latter problem is provided in Supplementary Materials D.
The issues with the density-ratio trick caused by different sample sizes could be addressed by applying different weights to the two groups so that the two groups have equal weights in total. Similar ideas of re-weighting the minor/major groups are proposed in the context of classification for unbalanced data, as, for example, the cost-sensitive boosting (Sun et al., 2007). To the best of our knowledge, the literature has not discussed either its application to density ratio estimation or, more importantly, its extension to Bayesian inference.
Connection 2: The Variational Form of the Squared Hellinger Distance
The design of the balancing loss can also be motivated from the perspective of estimating the an $f$-divergence between $p$ and $q$. In particular, in this section we show that the balancing loss turns out to be exactly equivalent to the variational form of one particular $f$-divergence, namely the squared Hellinger distance.
For a given convex function $f:[0,\infty)\to(-\infty,\infty]$ with $f(1)=0$, and $P$ is absolutely continuous with respect to $Q$, the general $f$-divergence is defined as $D_{f}(P\|Q):=\int f(\frac{p}{q})\,dQ$. It is known (Rényi, 1961; Csiszár, 1964; Ali and Silvey, 1966; Nguyen et al., 2010; Polyanskiy and Wu, 2025) that the dual form can be written as an optimization problem with respect to a functional class $\mathcal{F}$ mapping from $\mathcal{X}$ to $\mathbb{R}$,
$D_{f}(P\|Q)\geq\sup_{g\in\mathcal{F}}\int[g\,dP-f_{c}(g)\,dQ],$
where $f_{c}$ is the conjugate dual function associated with $f$, defined as $f_{c}(v):=\sup_{u\in\mathbb{R}}\{uv-f(u)\}$. The equality holds when the subdifferential $\partial f(q_{0}/p_{0})$ contains an element in $\mathcal{F}$. The subdifferential of $f$ is defined as $\partial f(t):=\left\{z\in\mathbb{R}:f(s)\geq f(t)+z(s-t),\forall s\in\mathbb{R}\right\}$. See details in Lemma 1 of Nguyen et al. (2010). The dual objective is referred to as the variational form of the $f$-divergence. In the case of the squared Hellinger divergence, $f(u)=\frac{1}{2}(\sqrt{u}-1)^{2}$, and the dual objective is $\sup_{g<\frac{1}{2}}\Big\{\mathbb{E}_{P}[g(X)]\;-\;\mathbb{E}_{Q}\!\Big[\frac{g(X)}{1-2g(X)}\Big]\Big\}$ with the optimum value $g^{*}=\frac{1}{2}-\frac{1}{2}\sqrt{\frac{q}{p}}$, based on a direct application of Lemma 1 in Nguyen et al. (2010). Now let $w=(1-2g)^{-1}$, the variational form for the squared Hellinger divergence becomes
$\sup_{w>0}\left\{1-\frac{1}{2}\int w^{-1}p\,d\mu-\frac{1}{2}\int wq\,d\mu\right\}$
which is clearly equivalent to minimizing the balancing loss. In other words, learning the density ratio by minimizing the balancing loss is essentially estimating the squared Hellinger distance using its variational form. Based on the known minimizer of the variational form, the minimizer of the balancing loss is $w^{*}=\sqrt{p/q}$. Hence, learning the density ratio can be achieved by minimizing the balancing loss.
Beyond these two additional motivations for the balancing loss, we will see later that the symmetry of the balancing loss, in $w$ vs $w^{-1}$ and $p$ vs $q$, will become handy when designing a generalized Bayesian framework to carry out inference on additive tree models.
2.3 Additive Tree Models and Boosting
Given the close relationship between density ratios and posterior odds, the latter of which can be estimated using supervised tree boosting with additive tree models, it is natural to consider the class of additive tree models as a promising candidate for approximating the density ratio. We thus consider from now on fitting additive tree models to the density ratio under the balancing loss. In this section, we begin with a non-Bayesian approach that aims to find a (functional) point estimate of the density ratio by optimizing the finite-sample version of the balancing loss. The resulting training algorithms are very similar to those used in supervised tree boosting.
Before introducing the algorithms, we describe the additive tree model. We consider a model for the log-balancing weight $\log w$ in the form
$\log w=\sum_{k=1}^{K}f_{k}$ (2)
where each base learner $f_{k}$ is a piecewise constant function defined on the leaf partition of the sample space given a tree $T_{k}$ with leaf node parameters $\beta_{k}=\{\beta_{k}(A)\}_{A\in T_{k}}$ and written as
$f_{k}=\sum_{A\in T_{k}}\beta_{k}(A)\mathbf{1}_{A}.$ (3)
Under this model, we are now presented with the following optimization problem:
$\min_{\{f_{k}\}}l_{n}(w)=\min_{\{\beta_{k},T_{k}\}}\frac{1}{n_{0}}\sum^{n_{0}}_{i=1}w^{-1}(x^{(0)}_{i})+\frac{1}{n_{1}}\sum^{n_{1}}_{i=1}w(x^{(1)}_{i})$
Next, we introduce two algorithms for fitting the above model under the balancing loss. These are direct counterparts of two different approaches to fitting additive trees in supervised learning problems. The first corresponds to a forward-stagewise (FS) algorithm based on greedy steepest descent on the space of single-tree-based piecewise-constant functions; the other corresponds to Friedman (2001)’s gradient boosting (GB) algorithm. In practice, these algorithms typically produce estimates that are very similar to each other, and so one could just adopt the simpler GB algorithm, but because the details of the FS algorithm shed more light on what constitutes effective function approximators for the density ratio under the balancing loss, we believe they are worthy of elaboration here.
Algorithm 1: A Forward-stagewise Algorithm
The following proposition characterizes the optimal single tree fit in each iteration of a forward-stagewise algorithm, which takes $K$ steps of iterative fitting and in each step $k$ finds the optimal fit of $f_{k}$ that minimizes the balancing loss. Its proof is provided in Supplementary Materials A.
Proposition 1.
Suppose $w_{k-1}$ is the estimate of $\sqrt{r^{*}}=\sqrt{p/q}$ after (k1)(k-1) steps of fitting. Then in the $k$th step, the loss is minimized by $\log w_{k}=\log w_{k-1}+f_{k}$ when the following conditions are satisfied by $f_{k}$:
  1. 1.
    Let $P_{k}|_{T_{k}}$ and $Q_{k}|_{T_{k}}$ be discrete distributions with their probability masses indexed by $A\in T_{k}$ such that
    $P_{k}|_{T_{k}}(A)\propto\int_{A}w_{k-1}^{-1}p\,d\mu,\ Q_{k}|_{T_{k}}(A)\propto\int_{A}w_{k-1}q\,d\mu.$
    Then $T_{k}$ is a partition structure that maximizes the Hellinger distance
    $H(P_{k}|_{T_{k}},Q_{k}|_{T_{k}})=\frac{1}{\sqrt{2}}\sqrt{\sum_{A\in T_{k}}\left(P_{k}|_{T_{k}}(A)-Q_{k}|_{T_{k}}(A)\right)^{2}.}$
  2. 2.
    For each $A\in T_{k}$, $\exp(\beta_{k}(A))=\sqrt{\int_{A}pw^{-1}_{k-1}\,d\mu}/\sqrt{\int_{A}qw_{k-1}\,d\mu}$.
Remark 1: If the balancing weight is correctly estimated, it achieves the symmetry between the two expectations in the loss in that $\mathbb{E}_{p}[w^{-1}]=\mathbb{E}_{q}[w]$, and we can impose this condition as a restriction or regularization on the estimator so that it does not focus on minimizing only one expectation out of the two. As justification, Proposition 1 implies that the balancing loss can be improved by multiplying the current estimate $w_{k}$ by a constant ($c$, say) such that $\mathbb{E}_{p}[(cw_{k})^{-1}]=\mathbb{E}_{q}[cw_{k}]$ (see Supplementary Materials A for details). Based on this discussion, we introduce this correction step in the boosting algorithms.
Remark 2: Given the connection between the balancing loss and the variational form of the Hellinger divergence, it is not surprising that the “optimal” single-tree in each iteration corresponds to the partition that maximizes the corresponding discretized approximation of the Hellinger divergence.
As in tree-based predictive models, searching for the exact optimal tree in each step is often not computationally tractable for multivariate data. Instead one can adopt a greedy, coarse-to-fine algorithm in which we iteratively choose a single cut point at each split of the tree $T_{k}$ that achieves the steepest ascent in the Hellinger distance or the steepest ascent in the corresponding affinity. Specifically, when we consider splitting the current node $A$, we can search for the splitting rule such that for the child nodes $A_{l}$ and $A_{r}$ the following Hellinger affinity is minimized, or, equivalently, the Hellinger distance is maximized:
$\sqrt{P_{k}|_{T_{k}}(A_{l})Q_{k}|_{T_{k}}(A_{l})}+\sqrt{P_{k}|_{T_{k}}(A_{r})Q_{k}|_{T_{k}}(A_{r})}.$
Algorithm 2: A Gradient Boosting Algorithm
We can also introduce a gradient boosting algorithm (Friedman, 2001), noting that the negative gradients of the empirical loss with respect to the additive estimand $F$ are as follows:
$\frac{\partial l_{n}(w)}{\partial F(x_{i}^{(0)})}=\frac{1}{n_{0}}\exp(-F(x_{i}^{(0)}))=\frac{1}{n_{0}}w^{-1}(x_{i}^{(0)}),\ -\frac{\partial l_{n}(w)}{\partial F(x_{i}^{(1)})}=\frac{1}{n_{1}}\exp(F(x_{i}^{(1)}))=-\frac{1}{n_{1}}w(x_{i}^{(1)}),$
which are the pseudo-residuals. One can then follow the recipe provided in (Friedman, 2001). Specifically, in each iteration of the algorithm we fit a regression tree to the pseudo-residuals by minimizing the squared error loss, that is, minimizing the variances of the gradients in the leaf nodes. We again adopt a coarse-to-fine greedy search and for each node $A$, we select the splitting rule with child nodes $A_{l}$ and $A_{r}$ that minimizes the variances. After constructing the tree, we obtain the optimal values of the node parameters given by Proposition 1.
Regularization strategies
As in supervised problems, the success of tree boosting under the two algorithms above requires each base learner to be regularized to extract only a small amount of structure in the density ratio function. We can regularize the contribution of each learner to the estimate by introducing a learning rate denoted by $\nu\in(0,1]$. If the current estimate is $w_{k-1}$ and the optimal learner obtained in the $k$th step is $f_{k}$, we can combine them as follows:
$\log w_{k}=\log w_{k-1}+\nu f_{k}.$
It is known that smaller values of the learning rate $\nu$ lead to better generalization performance in both supervised learning (Friedman, 2001; Hastie et al., 2009) and unsupervised learning (Awaya and Ma, 2024). Accordingly, in our implementation, we follow them by setting $\nu$ to, for example, $0.01$.
An additional, complementary strategy for regularization is to introduce a restriction on the size of trees to avoid constructing overly complicated partition structures, so we set the maximum resolution (depth) of the trees to be, for example, 4. Finally, the number of trees in the ensemble also needs to be carefully selected. We find that choosing this number via cross-validation yields satisfactory performance.
3 Generalized Bayesian Inference on Additive Trees
Tree boosting only offers point estimates for the density ratio function, without providing any associated uncertainty. To achieve uncertainty quantification, next we introduce a generalized Bayesian strategy (Bissiri et al., 2016) for learning the density ratio.
3.1 Balancing Loss as a Pseudo-likelihood
Following the framework introduced in (Bissiri et al., 2016), we define a pseudo-likelihood using the finite-sample loss $l_{n}(w)$, that is,
$L_{n,\tau}(w)=\tau^{2n_{\mathrm{min}}}\exp(-n_{\mathrm{min}}\tau l_{n}(w))$
where $n_{\min}$ is the sample size of the smaller group $\min(n_{0},n_{1})$, and $\tau>0$ is a “temperature” parameter. The term $\tau^{2n_{\mathrm{min}}}$ is added to provide a clearer interpretation of the pseudo-likelihood as a weighted likelihood under exponential distributions, as follows:
$L_{n,\tau}(w)=\prod^{n_{0}}_{i=1}\left[\tau\exp\left(-\tau w^{-1}(x^{(0)}_{i})\right)\right]^{\zeta_{0}}\times\prod^{n_{1}}_{i=1}\left[\tau\exp\left(-\tau w(x^{(1)}_{i})\right)\right]^{\zeta_{1}},$ (4)
where $\zeta_{0}=n_{\mathrm{min}}/n_{0}$ and $\zeta_{1}=n_{\mathrm{min}}/n_{1}$ are the weights taking values in $(0,1]$. If, for example, the first group has a smaller sample size with $n_{0}<n_{1}$, then the observations from the second group are down-weighted to keep the information from the two groups balanced. The effective sample size is controlled not by the total sample size but by $n_{\mathrm{min}}$, and this is natural since the amount of information about the two-sample difference provided by the data crucially depends on the size of the smaller group.
Under the formulation, the temperature $\tau$ can also be interpreted by a one-dimensional summary of the two-sample difference. To see this, we note that the likelihood $L_{n,\tau}(w)$ is maximized with respect to $\tau$ when $\tau^{-1}=\frac{1}{2}l_{n}(w)$. As $w$ is the balancing function corresponding to $\sqrt{p/q}$, what we estimate with $\tau^{-1}$ is
$\frac{1}{2}l_{n}(\sqrt{p/q})=\int\sqrt{pq}\,d\mu,$
which is the Bhattacharyya coefficient (Bhattacharyya, 1946).
Given the prior distribution of $w$, $\pi_{0}(w)$, the posterior distribution of $w$ is defined as
$\pi(w)\propto\pi_{0}(w)L_{n,\tau}(w),\$
where $\pi_{0}(w)$ is the prior of the balancing function $w$. Its specification is detailed next.
3.2 Conjugate Prior Specification
Next we describe prior specification on $w$ for efficient MCMC and a data-driven approach for setting the value of $\tau$. Recalling our model introduced in Eq. (2), we describe a Gibbs sampling strategy that iteratively samples from the full conditional of each basis function $f_{k}$, given all the others. That is, how one may update the node parameters $\beta_{k}(A)$ and the tree partition $T_{k}$ that define $f_{k}$ defined in Eq. (3).
3.2.1 Updating The Leaf Node Parameters
To derive the full conditional for $\beta_{k}(A)$, we define the balancing weight with $f_{k}$ subtracted:
$\log w_{-k}:=\log w-f_{k}.$
The portion of the likelihood $L_{n,\tau}(w)$ that depends on $\beta_{k}(A)$ is given by
$L_{n,\tau}(w)\propto\exp\left(-{\color[rgb]{0,0,0}{\zeta_{0}}}\tau\gamma_{k}(A)^{-1}\sum_{x^{(0)}_{i}\in A}w_{-k}^{-1}(x^{(0)}_{i})-{\color[rgb]{0,0,0}{\zeta_{1}}}\tau\gamma_{k}(A)\sum_{x^{(1)}_{i}\in A}w_{-k}(x^{(1)}_{i})\right)$,
where $\gamma_{k}(A)=\exp(\beta_{k}(A))$. Noting that this pseudo-likelihood resembles an exponential family kernel for $\gamma_{k}(A)$, we adopt an inverse-Gaussian leaf prior (Murray, 2021) for $\gamma_{k}(A)$, which is conjugate to this kernel. Specifically, the prior $\mathrm{InverseGauss}(\mu_{k}(A),\lambda_{k}(A))$ on $\gamma_{k}(A)$ is:
$p(\gamma_{k}(A)=z\mid\mu_{k}(A),\lambda_{k}(A))$ $=\sqrt{\frac{\lambda_{k}(A)}{2\pi z^{3}}}\exp\left(-\frac{\lambda_{k}(A)(z-\mu_{k}(A))^{2}}{2\mu_{k}(A)^{2}z}\right)$
$\propto\sqrt{\frac{\lambda_{k}(A)}{2\pi z^{3}}}\exp\left(-\frac{\lambda_{k}(A)}{2\mu_{k}(A)^{2}}z-\frac{\lambda_{k}(A)}{2z}\right).$
It is straightforward to see that the conditional posterior is analytically obtained in the form of $\mathrm{InverseGauss}(\mu^{\prime}_{k}(A),\lambda^{\prime}_{k}(A))$, where
$\mu^{\prime}_{k}(A)=\sqrt{\frac{\lambda_{k}(A)+2{\color[rgb]{0,0,0}{\zeta_{0}}}\tau\sum_{x^{(0)}_{i}\in A}w_{-k}^{-1}(x^{(0)}_{i})}{\frac{\lambda_{k}(A)}{\mu_{k}(A)^{2}}+2{\color[rgb]{0,0,0}{\zeta_{1}}}\tau\sum_{x^{(1)}_{i}\in A}w_{-k}(x^{(1)}_{i})}},\ \lambda^{\prime}_{k}(A)=\lambda_{k}(A)+2{\color[rgb]{0,0,0}{\zeta_{0}}}\tau\sum_{x^{(0)}_{i}\in A}w_{-k}^{-1}(x^{(0)}_{i})$.
We set $\mu_{k}(A)$ to $1$ so that the induced marginal prior on $\log w$ is around 0 (more discussions are to be given at the end of this section). The precision parameter $\lambda_{k}(A)$ induces shrinkage toward the prior mean 1, as seen in the expression of the posterior mean $\mu^{\prime}_{k}(A)$. We set $\lambda_{k}(A)$ according to the number of trees $K$, namely, $\lambda_{k}(A)=\lambda_{0}K,$ where $\lambda_{0}$ is a hyperparameter. The prior value of $\lambda_{k}(A)$ is linear in $K$ because the results in Whitmore and Yalovsky (1978) imply that the variance of $f_{k}(x)$ for each point $x$ ($x\in\Omega$) is $O(K^{-1})$ under our model specification (see Supplementary Materials A for more details). This choice of $\lambda_{k}(A)$ is also analogous to the amount of shrinkage recommended for the BART model (Chipman et al., 2010). We set $\lambda_{0}$ to 5 as a default value in all of our numerical experiments.
Another consideration is that the inverse Gaussian distribution has an asymmetric shape, so it can also introduce skewness into the prior of the log-balancing weight $\log w$. To avoid this issue, we apply the inverse Gaussian prior to $\gamma_{k}(A)$ and $\gamma_{k}(A)^{-1}$ alternately, that is,
$\gamma_{k}(A)\sim\mathrm{InverseGauss}(1,\lambda_{0}K)\ (k\text{: odd}),\ \gamma_{k}(A)^{-1}\sim\mathrm{InverseGauss}(1,\lambda_{0}K)\ (k\text{: even}).$
This ensures that the marginal prior of $\log w$ is exactly centered at 0 whenever the total number of trees $K$ is even.
3.2.2 Tree Priors and Posterior Updates on the Trees
We adopt the standard Bayesian CART prior (Chipman et al., 1998) on the tree $T_{k}$ as in BART (Chipman et al., 2010) and describe the prior on $T_{k}$ in terms of a recursive partition process. Starting from the full sample space as the root node, each node $A$ in the tree with depth $\mathrm{depth}(A)$ (which is 0 for the root) is further split with probability
$a_{T}(1+\mathrm{depth}(A))^{-b_{T}},\ a_{T}\in(0,1),b_{T}\in[0,\infty).$
We set the hyperparameters to the widely used default choices as $a_{T}=0.95$ and $b_{T}=2$ (Chipman et al., 2010). When splitting the node, one of the $d$ candidate dimensions of the sample space is selected with equal probability, and a cut point is selected uniformly over the range in that dimension. Following the standard recipe for BART (Chipman et al., 2010; Kapelner and Bleich, 2016; Sparapani et al., 2021), we update the tree $T_{k}$ by using Metropolis-Hastings moves. The details of the tree updating moves are provided in Supplementary Materials B.
3.3 Tuning the Temperature $\tau$
In generalized Bayesian inference, the temperature parameter $\tau$ controls the strength of the loss-based likelihood for the balancing weight function $w$ relative to the prior and significantly influences the resulting posteriors, including posterior variances and credible intervals. There is a growing body of literature on selecting the temperature parameter empirically, such as information matching (Holmes and Walker, 2017), calibrating frequentist coverage rates of credible intervals (Syring and Martin, 2019), and the SafeBayes algorithm (Grünwald and van Ommen, 2017). However, it is not straightforward to adapt these methods to the current context, where the task is to learn an unknown, nonparametrically specified function in an unsupervised setting and the computation cost of the MCMC is relatively expensive.
Instead, we adopt a hierarchical Bayesian approach by placing a prior on the temperature $\tau$, following Bissiri et al. (2016); Rigon et al. (2023). Motivated by the weighted likelihood expression in Eq. (4), we adopt a conjugate prior $\pi_{0}(\tau)$, namely $\mathrm{Gamma}(a_{0,\tau},b_{0,\tau})$, so the full conditional of $\tau$ is $\mathrm{Gamma}(a_{1,\tau},b_{1,\tau})$, where $a_{1,\tau}=a_{0,\tau}+2n_{\mathrm{min}},\ b_{1,\tau}=b_{0,\tau}+n_{\mathrm{min}}l_{n}(w).$
3.4 Empirical Validation of Uncertainty Calibration
We provide an illustration of the generalized Bayesian inference under two one-dimensional Gaussian distributions, $N(0,1)$ and $N(1,2.25)$, to verify whether the resulting generalized Bayesian uncertainty is properly calibrated.
The posterior distributions under the balanced and unbalanced sample sizes are provided in Figure 1, where the posterior means and the credible intervals are evaluated pointwise. We can see that the posterior distributions are centered around the true log-densities in a region with an adequate number of observations (i.e., the region between -2 and 2). Moreover, the inverse temperature $\tau^{-1}$, with which we estimate the Bhattacharyya coefficient, is distributed around the true value. We note that more information about the two-sample difference is provided under the balanced sample size as is represented by the weighted likelihood expression, and it is reflected by the different widths of the credible intervals in the posterior distributions. Figure 1 also shows the calibration of the credible intervals obtained by our Bayesian model in terms of nominal coverage rates based on 50 simulated data sets. It implies that the posterior uncertainty and the nominal coverage rate tend to agree more under a larger sample size, as in the case of $(n_{0},n_{1})=(2500,2500)$. The cases $(500,500)$ and $(500,4500)$ are similar, and it is consistent with the loss-based likelihood being driven by the smaller group size.
(a) Balanced ($n_{0}=500,n_{1}=500$)
(b) Unbalanced ($n_{0}=100,n_{1}=900$)
(c) Balanced ($n_{0}=2500,n_{1}=2500$)
(d) Unbalanced ($n_{0}=500,n_{1}=4500$)
Figure 1: An evaluation of posterior distributions under the 1D scenarios with the balanced/unbalanced sample sizes. Each left plot compares the true log-density ratios (gray, dotted) and the posteriors (mean: black, solid, 95% interval: blue, dashed) evaluated pointwise. Each middle plot shows the posterior distributions of the inverse temperature $\tau^{-1}$ and the true Bhattacharyya coefficient. Each right plot visualizes the calibration plot comparing the posterior uncertainty and the nominal coverage rate (namely, the ratio of observations included in the pointwise credible intervals) based on 50 simulations and their averages.
4 Numerical Experiments
We conduct simulation experiments to evaluate the proposed methods—both tree boosting and Bayesian additive trees. We begin with scenarios involving bivariate distributions, and then consider higher-dimensional (20-dimensional) cases, providing a comparison with several existing approaches. The first approach to be compared with is the “density-ratio trick” by inverting a binary classifier. In particular, we adopt AdaBoost for binary classification, as it uses the same class of additive tree ensembles to classify the two classes using essentially the exponential loss (Friedman et al., 2000). The difference between this approach and our boosting algorithms is essentially due to the different loss functions employed. To apply the density-ratio trick, we first estimate the classification probabilities and then transform them into the density ratios by multiplying the sample size ratios. For implementation, we use the ada function provided in the R package ada (Culp et al., 2016). In addition to the density-ratio trick, we evaluate the calibration method proposed in Cranmer et al. (2015) using AdaBoost as a discriminator. We also compare our methods with two popular kernel-based density ratio estimators, namely KLIEP (Sugiyama et al., 2008) and uLSIF (Kanamori et al., 2009), implemented with the R package densratio (Makiyama, 2019). For the proposed methods, we assess both types of boosting, the FS and the GB algorithms, along with the Bayesian additive tree model. Their implementation details are provided in Supplementary Materials C.
For each scenario, we evaluate the performance under two sample size settings: (i) $n_{0}=n_{1}=5000$ (the “balanced” setting) and (ii) $n_{0}=9000$, $n_{1}=1000$ (the “unbalanced” setting). As a measure of accuracy, we use the following symmetrized squared mean error for the true density ratio $r^{*}$ and the estimated ratio $\hat{r}$:
$\frac{1}{2n_{0}}\sum^{n_{0}}_{i=1}\left(\log r^{*}(x_{i}^{(0)})-\log\hat{r}(x_{i}^{(0)})\right)^{2}+\frac{1}{2n_{1}}\sum^{n_{1}}_{i=1}\left(\log r^{*}(x_{i}^{(1)})-\log\hat{r}(x_{i}^{(1)})\right)^{2},$
and for each setting we calculate the average error defined above based on 50 simulated datasets.
4.1 Two-dimensional Simulations
We simulate data sets using the following three scenarios which represent different types of distributional differences between $X_{0}$ and $X_{1}$ visualized in Figure 2. The first scenario represents a global-scale difference between the two samples, whereas the second and third scenarios involve more localized discrepancies. The details are provided in Supplementary Materials C
A comparison of the MSEs is provided in Table 1, and the results indicate that our proposed methods (the boosting and the Bayesian additive trees) generally show minimum errors. On the other hand, for the performance of the density ratio trick approach with AdaBoost, the MSEs increase substantially when comparing the balanced and unbalanced sample-size settings, reflecting the limitation discussed in Section 2.2. Also, the calibration method does not necessarily improve the performance, reflecting the fact that this approach involves a density estimation of the classification probabilities and thus can introduce an additional error. The two kernel-based methods (KLIEP and uLSIF) and the proposed methods based on the balancing loss are not strongly influenced by the unbalanced sample sizes, but the proposed method outperforms in all settings, due to the flexibility of the non-parametric models.
Figure 2: A representative sample of simulations generated from the three scenarios used for the two-dimensional experiments. The sample sizes, $n_{0}$ and $n_{1}$, are both set to 5000.
Table 1: A comparison of the MSEs and their standard errors in the two-dimensional scenarios.
Global Shift Local Shift Local Dispersion
$n_{0}/(n_{0}+n_{1})$ 0.5 0.9 0.5 0.9 0.5 0.9
GB 0.033 0.062 0.035 0.067 0.108 0.133
(0.001) (0.001) (0.001) (0.002) (0.002) (0.005)
FS 0.035 0.072 0.035 0.071 0.111 0.132
(0.001) (0.002) (0.001) (0.002) (0.002) (0.005)
BAT 0.010 0.018 0.046 0.138 0.112 0.190
(0.000) (0.001) (0.001) (0.002) (0.002) (0.006)
DRT (AdaBoost) 0.217 2.779 0.135 3.584 0.191 2.687
(0.022) (0.273) (0.011) (0.144) (0.010) (0.194)
CDC (AdaBoost) 0.117 0.533 0.789 2.099 1.286 2.731
(0.009) (0.048) (0.115) (0.280) (0.247) (0.415)
KLIEP 0.156 0.156 0.301 0.313 0.304 0.308
(0.005) (0.007) (0.004) (0.008) (0.004) (0.007)
uLSIF 1.068 0.422 0.383 0.410 0.142 0.211
(0.171) (0.096) (0.015) (0.022) (0.007) (0.010)
  • Notes: The methods are GB (gradient boosting), FS (forward stagewise), BAT (Bayesian additive trees), DRT and CDC (density-ratio trick and calibrated discriminative classifier based on AdaBoost), and the two kernel-based methods KLIEP and uLSIF.
Figure 3 compares the estimated log-density ratios in the local shift scenario with the balanced sample sizes ($n_{0}=n_{1}=5000$). The estimated functions using the proposed methods successfully capture the local difference structure, and importantly, the log-density ratios are significantly positive or negative for many observations included in the mixture component with the local shift difference. For example, in the plot of the 2.5% quantiles, we find observations with yellow/red colors, indicating that the log-density ratios are significantly larger than 0, mostly in the corresponding mixture component. Similar patterns are obtained for the other scenarios and sample size settings, as detailed in Supplementary Materials D.
Figure 3: The estimated log-density ratios obtained from the algorithms considered in the two-dimensional examples for the local shift scenario ($n_{0}=n_{1}=5000$). For Bayesian additive trees, the pointwise 2.5% quantiles and 97.5% quantiles of the log-density ratios are also displayed. The difference between FS and GB is minimal, so we show only the results for GB.
4.2 Higher-dimensional Settings
In this experiment, we evaluate the performance of the density ratio estimation algorithms in higher-dimensional settings. We generate 20-dimensional observations based on a latent factor model. First, we simulate latent variables $Z$ from 4-dimensional distributions, introducing the difference between the two samples in this latent space, and then project them onto the 20-dimensional space as
$\Lambda Z+\epsilon_{i},\ \epsilon_{i}\sim N(\mathbf{0},(0.1)^{2}I_{20}),$
where $\Lambda$ is a randomly generated orthonormal $20\times 4$ matrix satisfying $\Lambda^{\prime}\Lambda=I_{4}$. The distributions that the latent variables follow for the two groups, indicated by random variables $Z_{0}$ and $Z_{1}$, are defined under the two scenarios described below.
  1. 1.
    “Location Shift”: $Z_{0}\sim\frac{4}{5}N(\mu_{1},\Sigma_{1})+\frac{1}{5}N(\mu_{2},\Sigma_{2}),\ Z_{1}\sim\frac{4}{5}N(\mu_{1}+\delta,\Sigma_{1})+\frac{1}{5}N(\mu_{2},\Sigma_{2})$, where $\delta=(1,0,0,0)^{T}$.
  2. 2.
    “Dispersion Difference”: $Z_{0}\sim\frac{4}{5}N(\mu_{1},\Sigma_{1})+\frac{1}{5}N(\mu_{2},\Sigma_{2}),\ Z_{1}\sim\frac{4}{5}N(\mu_{1},\Delta\Sigma_{1})+\frac{1}{5}N(\mu_{2},\Sigma_{2})$, where $\Delta=\mathrm{diag}(0.5,1,1,1)$.
The parameters ($\mu_{k}$ and $\Sigma_{k}$ for $k=1,2$) are provided in Supplementary Materials C. The orthonormal matrix $\Lambda$ is generated using the randortho function in the R package pracma (Borchers, 2023).
Figure 4 shows the distribution of the first two latent factors in the first scenario, which account for the differences between the two groups, along with the corresponding true and estimated log-density ratios. The latent factor model for the true densities represents a challenging case for density ratio estimation, particularly for tree-based methods that use axis-aligned partitions along the original dimensions of the sample space. Nevertheless, as shown in Table 2, the proposed methods still produce more accurate estimates, and the performance is robust to the unbalanced sample sizes. More importantly, since estimating density ratios accurately becomes more challenging in such higher-dimensional settings, providing an uncertainty quantification for the estimates becomes even more crucial to conclude in which regions the two-sample difference exists. Our Bayesian approach serves this need by providing pointwise credible intervals for density ratios. The yellow/red regions in the lower quantiles and the blue regions in the upper quantiles correspond to the points at which the 95% credible bands exclude zero, indicating that the difference is significant at such points, and these indeed align with the regions with true differences.
Table 2: A comparison of the MSEs and their standard errors in the 20-dimensional scenarios.
Location Shift Dispersion
$n_{0}/(n_{0}+n_{1})$ 0.5 0.9 0.5 0.9
GB 0.073 0.125 0.151 0.227
(0.001) (0.002) (0.003) (0.005)
FS 0.076 0.134 0.158 0.237
(0.001) (0.002) (0.004) (0.005)
BAT 0.055 0.098 0.164 0.338
(0.001) (0.002) (0.004) (0.008)
DRT (AdaBoost) 0.230 3.718 0.150 3.103
(0.030) (0.152) (0.005) (0.171)
CDC (AdaBoost) 0.129 0.475 0.483 0.917
(0.007) (0.028) (0.025) (0.060)
KLIEP 0.614 0.607 0.547 0.548
(0.006) (0.006) (0.001) (0.001)
uLSIF 3.307 3.569 0.546 0.547
(0.070) (0.094) (0.001) (0.001)
  • Notes: The methods are GB (gradient boosting), FS (forward stagewise), BAT (Bayesian additive trees), DRT and CDC (density-ratio trick and calibrated discriminative classifier based on AdaBoost), and the two kernel-based methods KLIEP and uLSIF.
Figure 4: A comparison of the true log-density ratios and the estimated ratios projected onto the space of the latent variables $(z_{1},z_{2})$, for the location shift scenario ($n_{0}=n_{1}=5000$). For the Bayesian additive model, the 2.5% quantiles and 97.5% quantiles of the log-density ratios evaluated pointwise are also displayed. The difference in the results for FS and GB is minimal, so we present only the results for GB.
5 Assessing Synthesized Microbiome Compositions
In recent years, various generative models have been introduced to imitate sampling mechanisms of various complex biomedical data. In particular, several approaches have been proposed for generating microbiome compositions (Sankaran et al., 2025). A common need is to assess the quality of the synthetic samples drawn from these models by comparing them with the actual observed data. We apply our two-sample comparison framework to synthetic samples simulated from different generative models and the corresponding observed samples, thereby assessing the quality of data generation in these models.
The real microbiome data are obtained from the open-source Curated Metagenomic Data (Pasolli et al., 2017), available as the R package curatedMetagenomicData. We focus on the relative abundance data and apply the following subsetting rules: remove observations whose microbial relative abundance does not sum up to 1 (i.e., $\geq 0.99$); then, keep taxa with average relative abundance $\geq 0.1\%$; among those taxa, keep only observations that have zeros in fewer than 10% of taxa and whose mean abundance is at least the overall average. After subsetting, each row contains 123 columns of species-level relative abundance. We use one of the largest studies, the IBDMDB dataset (Lloyd-Price et al., 2019) in the database, as the data of interest. (The samples are in fact collected at multiple time points for each given subject, and the longitudinal nature of the data is important for scientific analysis. However, we ignore this aspect for the purpose of this demonstration since we will only focus on assessing how well generative models can approximate the marginal distribution of the samples.) We split the data into training ($n=1166$) and testing ($n=292$) sets. The generative models are fitted using the training data, and the testing data are used in the follow-up two-sample comparison with the synthetic samples. We evaluate various popular generative models, including both classical parametric models and state-of-the-art deep neural network-based models. The parametric models considered are the Dirichlet model (Holmes et al., 2012; La Rosa et al., 2012) and the Dirichlet tree model (Dennis III, 1991; Wang and Zhao, 2017; Mao and Ma, 2022). The nonparametric generators based on neural networks include continuous normalizing flows via conditional flow matching (Lipman et al., 2022; Tong et al., 2023, ICFM) (code available at https://github.com/atong01/conditional-flow-matching), and a generative adversarial network model called MB-GAN (Rong et al., 2021) (code available at https://github.com/zhanxw/MB-GAN), which is tailored for microbiome data generation. Our objective here is not to argue which model is better than others, as such judgments are context-dependent and require more careful tuning of each model. Rather, we aim to simply provide an example that density ratio-based two-sample comparison can provide insights into how various generative models can be gauged given the observations that they generate.
Train vs. Generated
Test vs. Generated
Figure 5: Principal Coordinates Analysis (PCoA) plots based on Bray-Curtis distance between observations. The blue and orange points represent the real sample and the generated sample, respectively. The first row presents the training sample vs. the generated sample, and the second row presents the test sample vs. the generated sample.
The PCoA plots based on the Bray-Curtis distance (Bray and Curtis, 1957) in Figure 5 provide a low-dimensional visualization for comparing the generated sample with the train/test sample. These plots clearly show that the parametric generators (Dirichlet and Dirichlet-Tree) have less coverage of the support than the train/test sample, while the two neural network-based approaches perform better in this regard. However, it is difficult to further differentiate the quality of the two neural network-based generators based on such visualization.
We apply the generalized Bayesian additive trees and obtain a posterior sample of the density ratio evaluated at each pair of the real and the synthetic samples. We then compute the posterior mean along with the pointwise 95% credible band for the density ratio, as shown in Figure 6(a) (training sample vs synthetic sample) and Figure 7(a) (testing sample vs synthetic sample). The color indicates the estimated (log) density ratio. Figure 6(b) and Figure 7(b) provide the PCoA visualization for the observation-level credible interval coverage, specifically, whether the credible interval for each sample is below 0, covering 0, or above 0. Figure 6(c) and 7(c) provide the credible intervals for the top 20 observations with the largest absolute log-density ratio values. Supplementary Materials D provides detailed plots on the credible bands for all samples. Recall that a log-density ratio of zero indicates that the generated and real samples are equally represented at the distributional level, which is the defining property of an ideal generative model. The estimated density ratios in both Figures 6 and 7 provide a clear contrast of the generation quality compared to the true sample. Compared with other generative models, the estimated log-density ratio for the MB-GAN-generated sample versus the real sample is closer to 0 across the support, especially in the test data. As shown in Figure 7(b), the credible bands provided by our proposed method offer uncertainty quantification for the density ratio, demonstrating 95% pointwise credible intervals of the log-density ratio include zero over most of the sample space for the MB-GAN, which is not observed for the parametric models and ICFM. If we compare the credible bands in Figures 6 and 7, because the training data size ($n=1166$) is much larger than the testing data ($n=292$), the posterior credible bands are generally narrower in the training data (Figure 6). This is especially clear for MB-GAN, where there are more observations with narrow credible bands that do not cover 0, meaning the pointwise posterior mass is more concentrated and away from 0; whereas in the testing data, almost all MB-GAN observations have credible intervals that cover 0.
(a) Train vs. Generated: Posterior mean of DRE
(b) Train vs. Generated: CI is below 0 (blue), contains 0 (grey), or above 0 (red).
(c) Observation-level credible interval for top 20 subjects with the largest absolute value of log-density ratio. The y-axis shows the sample index.
Figure 6: For the training data: (Left) PCoA plots (x-axis and y-axis are the first and second principal components) colored by the density ratio estimates of the training sample density over the generated sample density, and whether the 2.5% and 97.5% credible bands cover 0; (Right) For a few selected observations with the largest absolute log-density ratio averaged over all methods, the point estimates and the credible intervals for each generative model.
(a) Test vs. Generated: Posterior mean of DRE
(b) Test vs. Generated: CI is below 0 (blue), contains 0 (grey), or above 0 (red).
(c) Observation-level credible interval for top 20 subjects with the largest absolute value of log-density ratio. The y-axis shows the sample index.
Figure 7: For the test data: (Left) PCoA plots (x-axis and y-axis are the first and second principal components) colored by the density ratio estimates of the testing sample density over the generated sample density, and whether the 2.5% and 97.5% credible bands cover 0; (Right) For a few selected observations with the largest absolute log-density ratio averaged over all methods, the point estimates and the credible intervals for each generative model.
6 Conclusion
We have introduced a new approach to learning the density ratio between two i.i.d. samples using the class of additive tree models and a new loss function called the balancing loss. Not only does it provide effective boosting algorithms to attain functional (point) estimates, but it also provides a convenient means for attaining uncertainty quantification, which can be critical in challenging problems with limited sample size or high-dimensionality.
Additive tree models do have their limitations. In particular, they use axis-aligned partitions and therefore can sometimes be less effective at capturing high-order interactions, and they generally do not perform well with extremely high-dimensional data, i.e., those involving more than a few hundred dimensions. In such scenarios, an effective strategy is to couple the tree-based approaches with strategies for learning effective representations of the data. Examples of such representations might include low-dimensional subspace assumptions (e.g., manifolds) or other non-linear embeddings through state-of-the-art neural network-based encoders. Even in such cases, we believe that learning the density ratio between two samples is a worthy objective and will typically be easier to achieve than learning the density of each of the two samples per se.
Software
We have implemented an R package BATTS for the proposed boosting and Bayesian additive tree algorithms, and it is available at https://github.com/MaStatLab/BATTS .
Acknowledgment
This research is funded in part by NIGMS grant R01-GM135440 and NSF grant DMS-2152999. Part of the research was completed when LM and YX were at Duke University. NA’s research is supported by the Japan Society for the Promotion of Science KAKENHI (grant numbers: 25K16618, 24H00142, and 25H00546). Portions of the manuscript were prepared with the assistance of ChatGPT (from GPT-4.1 to 5.4) for English-language editing and coding.
Data Availability Statement
The real data, curated metagenomic data (Pasolli et al., 2017), used in Section 5 are openly available as an R package curatedMetagenomicData on GitHub and Bioconductor. The instruction and documentation are at https://waldronlab.io/curatedMetagenomicData/, DOI: 10.18129/B9.bioc.curatedMetagenomicData.
Appendix A Proofs
A.1 Proof of Proposition 1
We first fix the finite tree partition $T_{k}$. After adding the piecewise constant function $f_{k}$, the loss is given by
$l(w)=\sum_{A\in T_{k}}\left[\exp(-\beta_{k}(A))\int_{A}w^{-1}_{k-1}p\,d\mu+\exp(\beta_{k}(A))\int_{A}w_{k-1}q\,d\mu\right]$
By the relationship between the arithmetic mean and the geometric mean, each summand is bounded below by
$2\sqrt{\int_{A}w^{-1}_{k-1}p\,d\mu\int_{A}w_{k-1}q\,d\mu},$
and the equality is attained if
$\exp(-\beta_{k}(A))\int_{A}w^{-1}_{k-1}p\,d\mu=\exp(\beta_{k}(A))\int_{A}w_{k-1}q\,d\mu,$
and this condition yields the expression of the optimal $\beta_{k}(A)$. Furthermore, when the lower bound is attained for all $A\in T_{k}$, the loss can be expressed as
$l(w)=2\sum_{A\in T_{k}}\sqrt{\int_{A}w^{-1}_{k-1}p\,d\mu\int_{A}w_{k-1}q\,d\mu},$
and minimizing it with respect to $T_{k}$ is equivalent to maximizing the discrete Hellinger loss given in the claim. ∎
Also, we can derive the following corollary by considering the case in which the $T_{k}$ is the root tree.
Corollary 1.
Let $w$ be the current estimate of the balancing function and $c$ be the constant defined as $c=\sqrt{\mathbb{E}_{p}[w^{-1}]/\mathbb{E}_{q}[w]}$ and thus satisfies
$\mathbb{E}_{p}[(cw)^{-1}]=\mathbb{E}_{q}[(cw)].$
Then the balancing loss is improved as
$l(w)\geq l(cw).$
A.2 Derivation of the regularized prior for the node parameters
According to Whitmore and Yalovsky (1978), when $W\sim\mathrm{InverseGauss}(1,\theta)$ and $\theta$ are large, the distribution of the following variable is approximated by the standard normal distribution:
$\frac{1}{2\sqrt{\theta}}+\sqrt{\theta}\log W.$
Hence, when $\gamma_{k}(A)=\exp(\beta_{k}(A))$ follows $\mathrm{InverseGauss}(1,K\lambda_{0})$, the leaf node parameter $\beta_{k}(A)$ approximately follows
$N\left(-\frac{1}{2\sqrt{K\lambda_{0}}},\frac{1}{K\lambda_{0}}\right).$
This observation implies that, under the alternative prior
$\gamma_{k}(A)\sim\mathrm{InverseGauss}(1,K\lambda_{0})\ (k:\text{odd}),\ \gamma_{k}(A)^{-1}\sim\mathrm{InverseGauss}(1,K\lambda_{0})\ (k:\text{even}),$
the distribution of $\log w(x)$ for fixed $x$ is approximated by $N(0,\lambda_{0}^{-1})$.
Appendix B Details of the posterior sampling for the Bayesian additive tree model
B.1 MCMC algorithm
The Bayesian additive model includes three types of unknown variables: (i) tree partition structures $T_{1},\dots,T_{K}$, (ii) leaf node parameters $\beta_{1},\dots,\beta_{K}$, where $\beta_{k}$ denotes the collection of the parameters for the $k$th basis function, and (iii) the temperature $\tau$. Their posterior distributions are estimated with the following Metropolis-within-Gibbs algorithm:
  1. 1.
    For $k=1,\dots,K$,
    1. (a)
      Compute the residual $\log w_{-k}=\sum_{l\neq k}f_{l}$.
    2. (b)
      Update $T_{k}\mid w_{-k},\tau$.
    3. (c)
      Update $\beta_{k}\mid w_{-k},T_{k},\tau$.
  2. 2.
    Update $\tau\mid w$, where $w$ is the current balancing weight.
Updating the node parameters $\beta_{k}$ and the temperature $\tau$ is straightforward due to their conjugacy. The algorithm to update $T_{k}$ with the Metropolis-Hastings algorithm is detailed in the next section.
B.2 Updating trees
In this section, we consider updating the $k$th tree denoted by $T_{k}$ given the other parameter values using the proposals originally introduced for the Bayesian CART model (Chipman et al., 1998), and the details are obtained by following Chipman et al. (2010); Kapelner and Bleich (2016).
In each iteration of the MCMC, we propose a new tree structure $T^{*}_{K}$ by modifying the current tree $T_{k}$ via one of the three moves, GROW, PRUNE, and CHANGE, defined as follows:
GROW:
Randomly select one leaf node and split it to generate two child nodes.
PRUNE:
Randomly select one second-generation internal node (2GI), that is, a node with two child nodes that are terminal, and remove these child nodes.
CHANGE:
Randomly select one 2GI node and change the splitting rules by sampling a new rule from the prior distribution.
One of the moves is selected stochastically according to the fixed probabilities $p(\mathrm{GROW})$, $p(\mathrm{PRUNE})$, and $p(\mathrm{CHANGE})$, and they are all set to 1/3 in our numerical experiments.
Let $T^{*}_{k}$ denote the new proposed tree structure. Then, the proposal is accepted with probability $\min\{\alpha,1\}$, where
$\alpha=\underbrace{\frac{p(T^{*}_{k})}{p(T_{k})}}_{\text{Prior}}\underbrace{\frac{p(T^{*}_{k}\to T_{k})}{p(T_{k}\to T^{*}_{k})}}_{\text{Transition}}\underbrace{\frac{p(\mathrm{data}\mid T^{*}_{k},\{w_{-k}(x_{i})\}^{n}_{i=1})}{p(\mathrm{data}\mid T_{k},\{w_{-k}(x_{i})\}^{n}_{i=1})}}_{\text{Loss-based likelihood}}$
and $\{w_{-k}(x_{i})\}^{n}_{i=1}$ are the residuals. If a node $A$ is selected for updating $T_{k}$, then the ratios corresponding to the priors and the transition probabilities for the three moves are given as
[…]
(see Kapelner and Bleich (2016) for further details.)
For calculating the likelihood ratios, we need to note that we can analytically evaluate the likelihood with the leaf node parameters integrated out due to conjugacy as follows:
$\int p(\beta_{k})L_{n,\tau}(w)\,d\beta_{k}\propto\prod_{A\in T_{k}}L_{k,A},$
where $p(\beta_{k})$ is the joint prior density of the leaf node parameters, and
$L_{k,A}=\sqrt{\frac{\lambda_{k}(A)}{\lambda^{\prime}_{k}(A)}}\exp\left(\frac{\lambda_{k}(A)}{\mu_{k}(A)}-\frac{\lambda^{\prime}_{k}(A)}{\mu^{\prime}_{k}(A)}\right).$
Hence, for the GROW move in which the node $A$ is split into $A_{l}$ and $A_{r}$, the likelihood ratio is given as
$\frac{L_{k,A_{l}}L_{k,A_{r}}}{L_{k,A}},$
while in the PRUNE move, which removes $A_{l}$ and $A_{r}$, the ratio is the inverse. The CHANGE move generates two new children $A^{*}_{l}$ and $A^{*}_{r}$, so the likelihood ratio is given as
$\frac{L_{k,A^{*}_{l}}L_{k,A^{*}_{r}}}{L_{k,A_{l}}L_{k,A_{r}}}.$
Appendix C Details of experiments
C.1 2-dimensional numerical experiments
1. Global shift
$X_{0}\sim N\left(\mu^{(0)},I_{2}\right),\ X_{1}\sim N\left(\mu^{(1)},I_{2}\right)$, where $\mu^{(0)}=(-0.5,-0.5)^{\prime}$ and $\mu^{(1)}=(0.5,0.5)^{\prime}$.
2. Local shift
$X_{0}\sim\frac{1}{5}N(\mu_{1},\Sigma_{1})+\frac{1}{5}\sum^{5}_{k=2}N(\mu_{k},\Sigma_{k}),\ X_{1}\sim\frac{1}{5}N(\mu_{1}+\delta,\Sigma_{1})+\frac{1}{5}\sum^{5}_{k=2}N(\mu_{k},\Sigma_{k}),$
where $\delta=(0,1)^{\prime}$.
3. Local dispersion difference
$X_{0}\sim\frac{1}{3}N(\mu_{1},\Sigma_{1})+\frac{1}{3}\sum^{3}_{k=2}N(\mu_{k},\Sigma_{k}),\ X_{1}\sim\frac{1}{3}N(\mu_{1},\Delta\Sigma_{1})+\frac{1}{3}\sum^{3}_{k=2}N(\mu_{k},\Sigma_{k}),$
where $\Delta=\mathrm{diag}(0.36,1)$.
The parameters for simulating the data sets are specified as follows:
2. Local shift
$\mu_{1}=(9.0,9.9)^{\prime}$, $\mu_{2}=(-2.5,1.4)^{\prime}$, $\mu_{3}=(-2.3,-9.7)^{\prime}$, $\mu_{4}=(3.4,5.9)^{\prime}$, $\mu_{5}=(5.8,-9.5)^{\prime}$, $(\Sigma_{1}(1,1),\Sigma_{1}(1,2),\Sigma_{1}(2,2))=(2.9,0.5,1.1)$, $(\Sigma_{2}(1,1),\Sigma_{2}(1,2),\Sigma_{2}(2,2))=(1.2,-0.6,2.8)$,
$(\Sigma_{3}(1,1),\Sigma_{3}(1,2),\Sigma_{3}(2,2))=(2.3,-1.0,1.7)$, $(\Sigma_{4}(1,1),\Sigma_{4}(1,2),\Sigma_{4}(2,2))=(1.1,-0.4,2.9)$,
and $(\Sigma_{5}(1,1),\Sigma_{5}(1,2),\Sigma_{5}(2,2))=(3.0,0.2,1.0)$.
3. Local dispersion difference
$\mu_{1}=(1.9,-7.2)^{\prime}$, $\mu_{2}=(-2.3,-1.5)^{\prime}$, $\mu_{3}=(7.5,-3.1)^{\prime}$, $(\Sigma_{1}(1,1),\Sigma_{1}(1,2),\Sigma_{1}(2,2))=(1.0,-0.4,0.8)$, $(\Sigma_{2}(1,1),\Sigma_{2}(1,2),\Sigma_{2}(2,2))=(1.0,0,3.0)$, and $(\Sigma_{3}(1,1),\Sigma_{3}(1,2),\Sigma_{3}(2,2))=(2.9,0,1.1)$.
C.2 20-dimensional experiments
$\mu_{1}$ is set to $(-0.5,0,0,0)^{\prime}$ in the location shift scenario and $(0,0,0,0)^{\prime}$ in the dispersion difference scenario. The other parameters are fixed as $\mu_{2}=(0,0,0,0)^{\prime}$, $\Sigma_{1}=I_{4}$, and $\Sigma_{2}=4I_{4}$ in both scenarios.
C.3 Notes on the implementation of the density ratio estimators
For boosting, we assess both the FS and the GB algorithms. The learning rate $\nu$ is set to 0.01, and the maximum resolution of each tree is set to 4. The optimal number of trees is selected with the 5-fold cross-validation with the maximum number allowed set to 1000. In both types of the boosting algorithms, we introduce 31 equally spaced cut points in each dimension. For a fair comparison, AdaBoost is implemented under the same tuning parameters. For the generalized Bayesian additive tree model, we set the size of the ensemble to 200, following the standard implementation of the BART (Chipman et al., 2010). The MCMC sampler is initialized with trees containing only root nodes, and the size of the burn-in and the posterior sample is set to 2000 and 1000, respectively.
Additionally, in the boosting algorithms, we prune leaf nodes that contain observations from only one group to avoid division by zero in computing the optimal values of leaf node parameters.
For AdaBoost, we used the ada function from the ada package (Culp et al., 2016), and the optimal number of trees is selected via the $5$-fold cross-validation based on the classification error reported by the function. We found that the selected number tends to be extremely small when the two-sample difference is small (such as the local shift scenario in the 2D experiments), so we set the minimum number of trees to 10 manually. For the calibration method (Cranmer et al., 2015), we estimate the density functions of the classification probabilities for the two groups with the kernel density estimation implemented by the density function in R.
For the KLIEP and uLSIF, we use the functions KLIEP and uLSIF from the R package densratio (Makiyama, 2019) with the default parameter values. It should be noted that we detected an error in the source code for the KLIEP function  [^1] The estimated density ratios are not correctly computed on observations in a test data set in the original package. , so we used a fixed version that is available at https://github.com/nawaya040/densratio .
We found that the compared methods can output infinite values. Such values are omitted in computing the means and the standard errors of the results. We note that such phenomena were not observed for the proposed methods.
Appendix D Additional numerical experiment results
D.1 Comparison with the density ratio trick in the one-dimensional scenarios
We perform a simple experiment to discuss how the bias can be introduced by the density ratio trick based on classification methods, if the prior information on the sample size is not effectively incorporated into the learning algorithms.
As in Section 3.4, we simulate data sets from the two Gaussian distributions, $N(0,1)$ and $N(1,2.25)$, and compare two algorithms: (i) the density ratio trick based on the Real AdaBoost with ada function from the ada package Culp et al. (2016) and (ii) the gradient boosting algorithm that optimizes the proposed balancing loss function. For both algorithms, the maximum tree depth and the size of ensemble are set to 2 and 100, respectively, and the learning rate is 0.01. We set the sample size $(n_{0},n_{1})$ to $(500,500)$, $(300,700)$, $(200,800)$, and $(100,900)$, changing the degree of unbalancedness, and the results are shown in Figure 8. We can see that the estimated log-odds ratio from AdaBoost is overall shrunk toward 0, which is the initial value in the algorithm (Culp et al., 2007), and it results in the positive bias in estimating the density ratios under the unbalanced settings. In contrast, for our proposed approach, even though the scale of the log-density ratio is underestimated due to the finite sample, the sign of the estimated overall agrees with that of the truth. The same tendency of the AdaBoost-based method introducing biases is observed in the two-dimensional cases, as illustrated in Figure 10.
Figure 8: A comparison of the density-ratio trick (DRT) approach based on the AdaBoost and the proposed gradient boosting (GB) under the 1D example with different sample sizes. The first column shows a comparison between the true posterior log-odds ratio and those estimated by the AdaBoost, and the second column visualizes the log-density ratios estimated with DRT and GB along with the truth. Their estimation errors are visualized in the third column. The estimated functions are the averages of the results based on 50 simulated data sets.
D.2 Additional results for the two-dimensional scenarios
We provide additional visualizations corresponding to the settings described in Section 4.1. Figure 9 shows the estimated log-density ratios for the first (global shift) and third (local dispersion difference) scenarios under the balanced sample size setting, and Figure 10 shows the estimated ratios for the second (local shift) scenario under the unbalanced sample size configuration. For the unbalanced case, we note that the 95% credible intervals are above/below 0 for more than 10% of the observations, implying that the two-sample difference is significantly large in the corresponding regions, though this result is not clearly visible in the plots.
Figure 9: The estimated log-density ratios obtained using the algorithms considered in the two-dimensional examples for the global shift scenario and the local dispersion scenario ($n_{0}=n_{1}=5000$). For the Bayesian additive model, the 2.5% quantiles and 97.5% quantiles of the log-density ratios evaluated pointwise are also displayed. The difference between FS and GB is minimal, so we present only the results for GB.
Figure 10: The estimated log-density ratios obtained with the algorithms considered in the three scenarios with the unbalanced sample size ($n_{0}=9000$ and $n_{1}=1000$). For the Bayesian additive model, the 2.5% quantiles and 97.5% quantiles of the log-density ratios evaluated pointwise are also displayed. The difference between FS and GB is minimal, so we present only the results for GB.
We also show the calibration plots to evaluate the frequentist coverage rates of the credible intervals in Figure 11. We can observe only small over-coverage in the balanced scenarios, in which the data sets provide more information about the two-sample difference. In other words, we expect that the over-coverage and under-coverage under the unbalanced sample sizes are due to the structures of the differences that are difficult to accurately estimate with the small amount of information. For example, the diagonal shift in the global shift scenario is hard to describe with the axis-aligned tree partitions, and, for the local shift scenario, there is a sharp change in the density ratios in the local subset, but only a small number of observations from the second group (the smaller group) are observed there. Again, we note that such differences are accurately estimated under the balanced sample sizes in which richer information about the difference is provided from the data.
Figure 11: The calibration plots for the two-dimensional examples, in which the posterior uncertainty and the nominal coverage rates (namely, the ratio of observations included in the pointwise credible intervals) are compared under different probabilities. Each gray curve corresponds to one of the 50 simulated data sets, and the black curve indicates their averages.
D.3 Alternative simulation scenarios with 20-dimensional structures
In this section, we modify the simulation scenarios for the 20-dimensional cases so that the data structure becomes similar to the real microbiome data set analyzed in Section 5. To this end, we employ the following procedure:
  1. 1.
    We simulate a data set under the location shift scenario or the dispersion difference scenario.
  2. 2.
    We transform their marginal distributions by applying the CDFs of the Gaussian distributions  [^2] The parameters are selected so that the moments match those of the mixture component with the larger variances ($N(\mu_{2},\Sigma_{2})$). and then the inverse CDF of the beta distribution $\mathrm{Beta}(0.5,10)$.
As a result, the marginal distributions are right-skewed and have a large probability around 0, and the distributional structure is similar to the high-dimensional abundance data, as illustrated in Figure 12. The estimation results based on 50 simulated data sets are summarized in Table 3, and it shows that our proposed methods (GB, FS, and BAT) still estimate the two-sample difference more accurately.
Refer to caption Refer to caption
Figure 12: The first four marginal distributions of a simulated data set under the alternative 20-dimensional scenario based on “Location Shift”.
Table 3: A comparison of the MSEs and their standard errors in the alternative 20-dimensional scenarios.
Location Shift Dispersion
$n_{0}/(n_{0}+n_{1})$ 0.5 0.9 0.5 0.9
GB 0.077 0.129 0.163 0.236
(0.001) (0.002) (0.004) (0.005)
FS 0.080 0.138 0.171 0.248
(0.001) (0.002) (0.005) (0.006)
BAT 0.062 0.107 0.179 0.361
(0.001) (0.002) (0.004) (0.008)
DRT (AdaBoost) 0.243 3.651 0.149 2.940
(0.030) (0.152) (0.004) (0.163)
CDC (AdaBoost) 0.133 0.495 0.469 0.896
(0.008) (0.043) (0.024) (0.047)
KLIEP 0.682 0.681 0.538 0.539
(0.003) (0.004) (0.001) (0.001)
uLSIF 19.278 11.592 0.538 0.539
(7.435) (5.744) (0.001) (0.001)
  • Notes: The methods are GB (gradient boosting), FS (forward stagewise), BAT (Bayesian additive trees), DRT and CDC (density-ratio trick and calibrated discriminative classifier based on AdaBoost), and the two kernel-based methods KLIEP and uLSIF.
Appendix E Additional real data analysis results
Figure 13 and 14 provide the additional real data analysis results on the credible interval comparison between different generative models. The credible bands for all samples are shown in the PCoA plots.
Train vs. Generated: Posterior mean of DRE
Train vs. Generated: Lower 2.5% MCMC quantile of DRE
Train vs. Generated: Upper 97.5% MCMC quantile of DRE
Figure 13: PCoA plots colored by the density ratio estimates of the training sample density over the generated sample density, along with the 2.5% and 97.5% credible bands.
Test vs. Generated: Posterior mean of DRE
Test vs. Generated: Lower 2.5% MCMC quantile of DRE
Test vs. Generated: Upper 97.5% MCMC quantile of DRE
Figure 14: PCoA plots colored by the density ratio estimates of the testing sample density over the generated sample density, and the 2.5% and 97.5% credible bands.